### v2

# From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles

Abstract
Model combination is a powerful approach for achieving superior performance compared to selecting a single model. We study both theoretically and empirically the effectiveness of ensembles of Multi-Frequency Echo State Networks (MFESNs), which have been shown to achieve state-of-the-art macroeconomic time series forecasting results (Ballarin et al., 2024a, ). The Hedge and Follow-the-Leader schemes are discussed, and their online learning guarantees are extended to settings with dependent data. In empirical applications, the proposed Ensemble Echo State Networks demonstrate significantly improved predictive performance relative to individual MFESN models.
111*
YCL and LG acknowledge the financial support of the FoKo of the University of St. Gallen (Project Nr. 1022324, “Machine Learning Techniques for Macroeconomic Forecasting and Pricing of Financial Derivatives”). GB thanks the Great Minds Postdoctoral Fellowship program of the University of St. Gallen, which made this research possible. Python code accompanying this manuscript is available at the online repository https://github.com/RCEconModelling/Ensemble-MFESN
Keywords: echo state networks, model combination, online learning, time series prediction
1 Introduction
Many mathematical, statistical, and econometric tools have been developed to solve the problem of predicting stochastic and deterministic processes. In empirical applications, however, it is often difficult to identify which model, or even which model class, is most appropriate for a given task. Models that perform well on average may fail during particular periods. This situation is common in forecasting settings, for example, involving macroeconomic time series. Structural breaks, financial crises, and policy regime shifts can significantly affect predictive accuracy. Since forecasting performance can vary substantially over time, it is natural, when model updating is infeasible or unsuitable, to revise the choice of model(s) as new information becomes available.
The model combination approach formalizes the use of multiple prediction schemes. The forecaster relies on a finite set, referred to as an ensemble, of distinct models or algorithms, termed experts, to predict an outcome of interest. Experts in an ensemble are evaluated and reweighted over time based on their predictive losses. Ensemble methods can hence adapt to changing time series dynamics by favoring different models across periods. The objective is to construct data-driven weights that aggregate individual expert forecasts to minimize cumulative loss over time. A well-designed combination scheme is expected to select the best-performing models, achieving performance close to the minimum loss in hindsight.
Forecast combination is a long-standing idea in econometrics and time series analysis (Bates and Granger,, 1969; Winkler and Makridakis,, 1983; Clemen,, 1989; Granger,, 1989) [^2] Boosting, bagging, and random forest methods, which explicitly combine many learners into one predictor to improve performance, have also become increasingly popular (Athey,, 2019; Athey and Imbens,, 2019)., with a large literature documenting its empirical effectiveness in a variety of macroeconomic and financial forecasting contexts. Subsequent contributions have studied both linear and nonlinear combination schemes, as well as the role of time variation in forecast weights (see Timmermann,, 2006, for a comprehensive survey). More recent work has emphasized adaptive and data-driven approaches to forecast combination in high-dimensional and potentially unstable environments, including settings with structural change and model misspecification (e.g., Hansen, (2008); Cheng and Hansen, (2015); Kim and Swanson, (2018); den Reijer and Johansson, (2019); Bolhuis and Rayner, (2020); Fuleky, (2020); Masini et al., (2023)), as well as considered Bayesian (Wang et al.,, 2023) and density combination schemes (Wallis,, 2011).
The literature on online combination schemes and ensembles is rich. In the standard framework of prediction with expert advice (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997; Vovk,, 1995; Cesa-Bianchi and Lugosi,, 2006), or its variant known as the Hedge setting, much of the work focuses on developing combination algorithms that enjoy guarantees on regret, defined as the difference between the cumulative loss of the forecaster and that of the best among $K$ experts. Prominent examples include Follow-the-Leader (FTL), which assigns uniform weights at each round to the model(s) with the smallest cumulative loss so far; it is known to achieve an $O(\log T)$ regret for strongly convex losses in stochastic settings (with $T$ a time horizon), while potentially performing poorly under worst-case data (van Erven et al.,, 2011). The Hedge algorithm (Freund and Schapire,, 1997), also called the exponentially weighted forecaster, provides a smoothed alternative to FTL and enjoys an anytime worst-case $O(\sqrt{T\log K})$ regret bound under an optimal learning rate. In stochastic environments, allowing the learning rate of Hedge to be data-adaptive yields regret comparable to FTL, motivating schemes such as Adaptive Hedge (van Erven et al.,, 2011; de Rooij et al.,, 2014). Significant effort has been devoted to develop strategies that combine the best properties of these approaches (see, for example, Koolen et al., (2014); Wintenberger, (2017); Mourtada and Gaïffas, (2019); Ito et al., (2024); Wintenberger, (2024)).
In this paper, we bridge the literature on online ensemble learning (see, e.g., Shalev-Shwartz,, 2012) with econometric forecast combination (Timmermann,, 2006) in a general time series setup. From a theoretical perspective, we study the properties of Follow-the-Leader and Decreasing Hedge (DecHedge) algorithms in the stochastic setting for i.i.d. (in time) and $\varphi$-mixing losses, building upon the developments in Mourtada and Gaïffas, (2019) and de Rooij et al., (2014), respectively. For both algorithms, we derive Hoeffding- and Bernstein-type regret bounds that characterize their cumulative performance relative to the best expert in hindsight. While the dependence on the number of experts is mild (logarithmic), the gap between the least expected loss and the runner-up plays a key role in our bounds.
On the empirical side, we construct our ensembles using Multi-Frequency Echo State Network (MFESN) models introduced in Ballarin et al., 2024a , which have been shown to outperform state-of-the-art mixed-frequency forecasting methods such as MIxed DAta Sampling (MIDAS) (Ghysels et al.,, 2007) and Dynamic Factor Models (DFMs) (Geweke,, 1977; Sargent and Sims,, 1977). The MFESN models are based on Echo State Networks (ESNs), a class of recurrent-type neural networks with randomly sampled state parameters (Jaeger,, 2010; Maass,, 2011; Lukoševičius and Jaeger,, 2009). This model family has been actively studied in terms of approximation, generalization, and memory properties (Grigoryeva and Ortega, 2018b, ; Grigoryeva and Ortega, 2018a, ; Gonon and Ortega,, 2020; Gonon et al., 2020b, ; Gonon et al.,, 2023; Ballarin et al., 2024b, ). Using quarterly U.S. GDP growth and a set of $33$ monthly and daily financial and macroeconomic variables in Ballarin et al., 2024a , we show that MFESN ensembles substantially reduce mean-squared forecasting errors. Additionally, allowing for heterogeneity in random parameter draws and state leak rates across the ensemble yields up to $40\%$ improvements relative to baseline MFESNs. These results set new benchmarks in the prediction of U.S. GDP growth with multi-frequency regressors. Finally, we report the experts’ weight evolution in our forecasting exercise, showing that a handful of models receive most of the weight.
[…]
Notation.
[…]
2 Preliminaries
In this section, we introduce the general econometric setting of interest, which involves constructing time series predictions by combining (or selecting) models from a set of experts.
2.1 General Setup
Let $\{Y_{t}\}_{t\in\mathbb{Z}}$, $Y_{t}\in\mathbb{R}$, be the target time series, and $\{Z_{t}\}_{t\in\mathbb{Z}}$, $Z_{t}\in\mathbb{R}^{d}$, a vector of regressors, which may include lagged values of $Y_{t}$. We consider the problem of predicting the target series $Y_{t+h}$ at the forecasting horizon $h\in\mathbb{N}^{+}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a fixed probability space on which all the random variables are defined. For each $t\in[T]$, let $Z_{t}$ be $\mathcal{F}_{t}$-measurable with $\mathcal{F}_{t}:=\sigma\left(Z_{t}\right)$. Under relatively weak conditions of the joint law of $Y_{t}$ and $Z_{t}$ the $h$-step-ahead minimum mean square error (MSE) predictor of $Y_{t+h}$ at time $t$, with $h\in[H]$, is the conditional expectation function (CEF) given by
$\mathbb{E}[Y_{t+h}\,|\,Z_{t}]:=\arg\inf_{g\in\mathcal{G}}\mathbb{E}\left[\|Y_{t+h}-g(Z_{t})\|_{2}^{2}\>\big|\,\mathcal{F}_{t}\right],$ (1)
where $\mathcal{G}$ is the class of $\mathcal{F}_{t}$-measurable functions.
The core challenge arising when solving (1), even in the one-step-ahead prediction setting ($h=1$), is that minimizing MSE over (possibly infinite-dimensional) $\mathcal{G}$ class of functions can be infeasible or computationally and statistically demanding. This is typically addressed by structural assumptions on the joint data-generating process $(Y_{t},Z_{t})$. For example, standard choices include linear models (Brockwell and Davis,, 2006), semi- or nonparametric classes of sufficient regularity (Stone,, 1982; Li and Racine,, 2009; Tsybakov,, 2009), or a reproducing kernel Hilbert space setting (Christmann and Steinwart,, 2008; Schölkopf et al.,, 2002; Berlinet and Thomas-Agnan,, 2011). In this work, we study a standard online aggregation setting in which multiple models (“experts”) are combined, each potentially excelling under different finite sample realizations, but always assuming that a best-performing model exists (see Section 3). The focus of the paper is on studying schemes to weight experts in order to produce a combined forecast that over time aligns with the optimal expert.
2.2 Prediction with Experts
The concept of prediction with experts goes back to the works of Littlestone and Warmuth, (1994); Foster, (1991); Foster and Vohra, (1993); Vovk, (1995); Cesa-Bianchi et al., (1997). A forecaster is interested in obtaining, at each time step $t\in[T]$, a prediction $\widehat{Y}_{t+h}$, for a horizon $h\in[H]$, based on $K\geq 1$ experts indexed by $k\in[K]$, each providing their own forecast $\widehat{Y}_{t+h}^{(k)}$. The forecaster constructs a convex combination of $K$ expert predictions using a weight vector $\bm{\omega}_{t,h}:=(\omega^{(1)}_{t,h},\ldots,\omega^{(K)}_{t,h})^{\top}\in\mathbb{R}^{K}$ such that
$\bm{\omega}_{t,h}\geq 0\quad\text{and}\quad\bm{\omega}_{t,h}^{\top}\bm{1}_{K}=1\kern 5.0pt\text{ for all }t\in[T]\text{ and for all }h\in[H],$
and producing a forecast
$\widehat{Y}_{t+h}:=\sum_{k=1}^{K}\omega_{t,h}^{(k)}\widehat{Y}_{t+h}^{(k)}.$ (2)
We will also use $\Delta^{K-1}:=\{\bm{\omega}\in\mathbb{R}_{+}^{K}:\bm{\omega}^{\top}\bm{1}_{K}=1\}$ to indicate the real simplex.
The accuracy of each forecast is assessed using a loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$. We define, for each expert $k\in[K]$ and each $t\in[T]$, the associated loss at $h$-horizon forecasting task of predicting ${Y}_{t+h}$, $h\in[H]$, as
$\ell^{(k)}_{t,h}:=\ell({Y}_{t+h},\widehat{Y}_{t+h}^{(k)}),$
the vector containing the losses of all $K$ experts as $\bm{\ell}_{t,h}:=(\ell^{(1)}_{t,h},\ldots,\ell^{(K)}_{t,h})^{\top}$ and the forecaster’s loss $\overline{\ell}_{t,h}:=\bm{\omega}_{t,h}^{\top}\bm{\ell}_{t,h}$. We also define the cumulative loss of expert $k$ and of the forecaster, incurred by time $t$ when forecasting $h$ steps ahead, as
$L^{(k)}_{t,h}:=\sum_{\tau=1}^{t-h}\ell^{(k)}_{\tau,h},\quad{\rm and}\quad\overline{L}_{t,h}:=\sum_{\tau=1}^{t-h}\overline{\ell}_{\tau,h},$ (3)
respectively, with the convention that $L^{(k)}_{t,h}=0$ and $\overline{L}_{t,h}=0$ for all $t\leq h$, and construct a vector $\bm{L}_{t,h}:=(L^{(1)}_{t,h},\ldots,L^{(K)}_{t,h})^{\top}$ of cumulative losses of all $K$ experts.
The natural objective of online learning procedures is to minimize the loss $\overline{L}_{t,h}$ experienced by the forecaster when combining $K$ experts’ predictions. Assuming that $[K]$ does not change over time, one can introduce cumulative regret at $t\in[T]$ for horizon $h\in[H]$ as
$\overline{R}_{t,h}:=\overline{L}_{t,h}-\min_{k\in[K]}L^{(k)}_{t,h}$, (4)
[…]
Throughout the paper, unless stated otherwise, we impose the following assumptions on expert and forecaster losses for all $h\in[H]$.
Assumption 1 (Bounded losses).
Losses $\ell^{(k)}_{t,h}\in[0,1]$ almost surely for all $k\in[K]$, $t\in\mathbb{N}$.
Assumption 2 (Temporal dependence).
Losses $\{\bm{\ell}_{t,h}\}_{t}$, $\bm{\ell}_{t,h}\in\mathbb{R}^{K}$, satisfy one of the following:
$\{\ell^{(k)}_{t,h}\}_{t}$ is i.i.d. for every $k\in[K]$;
$\{\ell^{(k)}_{t,h}\}_{t}$, for every $k\in[K]$, is strictly stationary and $\varphi$-mixing with coefficients $\{\varphi^{(k)}_{n}\}_{n\geq 1}$ depending only on the lag $n$;
The vector-valued process $\{\bm{\ell}_{t,h}\}_{t}$ is strictly stationary and $\varphi$-mixing with coefficients $\{\varphi_{n}\}_{n\geq 1}$ depending only on the lag $n$.
The definition of $\varphi$-mixing dependence in (ii) and (iii) is in the sense of Definition A.1 in the Appendix, following Rio, (2017). The dependence conditions are imposed on losses (hence, data and experts) over time, but no independence is assumed across the $K$ losses.
Remark 2.1.
Under Assumption 1 and Assumption 2, for all $k\in[K]$, $\{\ell^{(k)}_{t,h}\}_{t}$ is a stationary process with time-invariant mean $\mu_{k,h}=\mathbb{E}[\ell^{(k)}_{t,h}$] and variance $v_{k,h}={\rm Var}[\ell^{(k)}_{t,h}$].
Assumption 3 (Uniform summability).
Losses $\{\bm{\ell}_{t,h}\}_{t}$, $\bm{\ell}_{t,h}\in\mathbb{R}^{K}$, satisfy one of the following:
There exists a finite constant $C_{1,\varphi}>0$ such that $\max_{k\in[K]}\sum_{n=1}^{\infty}\varphi_{n}^{(k)}<C_{1,\varphi},$ with $\{\varphi^{(k)}_{n}\}_{n\geq 1}$ coefficients of $\{\ell^{(k)}_{t,h}\}_{t}$, for every $k\in[K]$.
There exists a finite constant $\widetilde{C}_{1,\varphi}>0$ such that $\sum_{n=1}^{\infty}\varphi_{n}<\widetilde{C}_{1,\varphi}$ with $\{\varphi_{n}\}_{n\geq 1}$ coefficients of vector-valued $\{\bm{\ell}_{t,h}\}_{t}$.
Assumption 4 (Uniform root-summability).
Losses $\{\bm{\ell}_{t,h}\}_{t}$, $\bm{\ell}_{t,h}\in\mathbb{R}^{K}$, satisfy one of the following:
There exists a finite constant $C_{2,\varphi}>0$ such that $\max_{k\in[K]}\sum_{n=1}^{\infty}\sqrt{\varphi_{n}^{(k)}}<C_{2,\varphi},$ with $\{\varphi^{(k)}_{n}\}_{n\geq 1}$ coefficients of $\{\ell^{(k)}_{t,h}\}_{t}$, for every $k\in[K]$.
There exists a finite constant $\widetilde{C}_{2,\varphi}>0$ such that $\sum_{n=1}^{\infty}\sqrt{\varphi_{n}}<\widetilde{C}_{2,\varphi}$ with $\{\varphi_{n}\}_{n\geq 1}$ coefficients of vector-valued $\{\bm{\ell}_{t,h}\}_{t}$.
Assumption 1 is standard in online learning and, up to a fixed rescaling, requires losses to be bounded. Assumption 2 imposes conditions on the temporal dependence of loss vectors, as the prediction algorithm evolves. While much of the literature focuses on the i.i.d. case, mixing conditions such as Assumption 2(ii) are empirically more relevant in time-series applications. Notice also that for any $k\in[K]$, since $0\leq\varphi_{n}^{(k)}\leq 1$, Assumption 4(i) implies Assumption 3(i) and Assumption 4(ii) implies Assumption 3(ii). Moreover, in each of these assumptions, part (ii) implies part (i). All assumptions are imposed on the loss processes $\{\ell^{(k)}_{t,h}\}_{t}$, rather than on the joint data process $(Y_{t},Z_{t})$. If experts are fixed and $(Y_{t},Z_{t})$ is mixing, the induced loss processes inherit this property under measurable maps. Even under mild nonstationarity (e.g., local stationarity), these losses often remain weakly dependent, making uniform mixing a plausible assumption in practice.
In the remainder of the paper, we focus on a single forecasting horizon $h\in\mathbb{N}$ and hence drop the $h$ subscript where convenient. Our results extend directly to the multi-horizon case by analyzing cumulative regret separately for each horizon $h\in[H]$.
3 Expert Ensembles
We now turn to an in-depth discussion of the Follow-the-Leader and Hedge online learning methods. Table 1 provides an overview of these schemes, as well as classical combination approaches based on averaging and rolling mean-squared-error weighting (see Appendix D for more details). In Section 3.3, we theoretically study the stochastic setting with and without temporal dependence, and derive regret guarantees for both FTL and decreasing Hedge.
Type Scheme Weighting, $\omega^{(k)}_{t}$ Specification Reference
Direct Simple Averaging $\frac{1}{K}$ Bates and Granger, (1969) Timmermann, (2006)
Rolling MSE $\frac{\left({\mathrm{MSE}}^{(k)}_{t,r}+\varepsilon\right)^{-1}}{\sum_{j=1}^{K}\left({\mathrm{MSE}}^{(j)}_{t,r}+\varepsilon\right)^{-1}}$ fixed constant $\varepsilon>0$ rolling window length $r$
Follow-the-Leader $\frac{1}{|\textnormal{{ftl}}_{t}|}\mathds{1}\{k\in\textnormal{{ftl}}_{t}\}$ $\textnormal{{ftl}}_{t}=\arg\min_{j\in[K]}L^{(j)}_{t-1}$ de Rooij et al., (2014)
Hedge Constant $\frac{\exp\left(-\eta_{t}L^{(k)}_{t-1}\right)}{\sum_{j=1}^{K}\exp\left(-\eta_{t}L^{(j)}_{t-1}\right)}$ with learning rate $\eta_{t}$ $\eta_{t}=\eta\propto\sqrt{\frac{\log K}{T}}$ Littlestone and Warmuth, (1994) Cesa-Bianchi and Lugosi, (2006) Chernov and Zhdanov, (2010)
Doubling $\eta_{t}\propto\sqrt{\frac{\log K}{2^{r-1}}}$ doubling round $r=\left\lceil\log_{2}(t+1)\right\rceil$
Decreasing $\eta_{t}\propto\sqrt{\frac{\log K}{t}}$
Adaptive $\eta_{t}=\frac{\log K}{\nabla_{t-1}}$ cumulative mixability gap $\nabla_{t-1}$ van Erven et al., (2011) de Rooij et al., (2014)
Table 1: Overview of online ensemble learning and combination schemes considered.
3.1 Follow-the-Leader
An intuitive first step to construct expert weights is the so-called Follow-the-Leader (FTL) strategy: At each round, uniform weights are assigned to the (subset of) model(s) with the smallest cumulative loss so far. Formally, let $\textnormal{{ftl}}_{t}\in\arg\min_{k\in[K]}L^{(k)}_{t-1}$ be the subset of leader-experts at time $t$, that is, the experts with the least cumulative loss up to t1t-1. The Follow-the-Leader weights are defined as
$\omega^{(k)}_{\textnormal{{ftl}},t}=\frac{1}{|\textnormal{{ftl}}_{t}|}\mathds{1}\{k\in\textnormal{{ftl}}_{t}\},$ (5)
with the convention that $\omega^{(k)}_{\textnormal{{ftl}},1}=1/K$. FTL is known to perform well in several settings, particularly in the case of i.i.d. losses. However, it has the downside of being sensitive to adversarially-generated losses, as it overreacts to short-term fluctuations. de Rooij et al., (2014) derive an upper bound on the cumulative regret incurred by a forecaster using the FTL strategy due to the variability of the losses and the number of leader changes.
Let $s_{t}:=\max_{k\in[K]}\ell^{(k)}_{t}-\min_{k\in[K]}\ell^{(k)}_{t}$ and $S_{t}:=\max\{s_{1},\ldots,s_{t}\}$ denote the instantaneous maximal loss differential and maximum loss differential over $t$ periods, respectively. Define $c_{t}=1$ if there exists an expert $k^{\prime}$ with $k^{\prime}\in\mathrm{FTL}_{t-1}$ but $k^{\prime}\notin\mathrm{FTL}_{t}$ (i.e., the leader set changes at $t$), and $c_{t}=0$ otherwise. We can write $c_{t}=\mathds{1}\{\textnormal{{ftl}}_{t-1}\not=\textnormal{{ftl}}_{t}\}$. Let then $C_{t}:=\sum_{\tau=1}^{t}c_{\tau}$ be the total number of times the leader set changes up to time $t\in[T]$.
Lemma 3.1 (de Rooij et al., 2014, Lemma 10).
The cumulative FTL regret satisfies
$\overline{R}_{\textnormal{{ftl}},T}\leq S_{T}C_{T}.$
Lemma 3.1 shows that FTL is effective whenever leader changes are infrequent, and the loss range is small. Under i.i.d. losses whose expectations are well separated, frequent leader changes are unlikely (de Rooij et al.,, 2014). In Section 3.3, we provide a sharper result in a stochastic setting, while also allowing for dependence.
3.2 Hedge and Adaptive Hedge
Even though FTL relies on identifying a single best expert for prediction, the underlying setup can be generalized to non-sparse weighting vectors. The overarching framework is that of the Hedge family of algorithms (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997, 1999; Cesa-Bianchi et al.,, 1997), also known as exponentially weighted average forecasters. In Hedge, experts are weighted according to
$\omega^{(k)}_{\textnormal{{hdg}},t}=\frac{\exp(-\eta_{t}L^{(k)}_{t-1})}{\sum_{j=1}^{K}\exp(-\eta_{t}L^{(j)}_{t-1})}$, (6)
where $\{\eta_{t}\}_{t=1}^{T}$, with $\eta_{t}>0$ for all $t\in[T]$, is the sequence of the so-called learning rates (Cesa-Bianchi and Lugosi,, 2006). We use the standard convention $\omega^{(k)}_{\textnormal{{HDG}},1}=1/K$. Variations of Hedge propose different sequences of learning rates aimed at minimizing the cumulative regret of the forecaster. A natural way to interpret (6) is to note that it corresponds to the gradient of a scaled LogSumExp function of the negative cumulative losses
$\omega^{(k)}_{\textnormal{{hdg}},t}=\frac{\partial\mathcal{L}(\bm{L}_{t-1},\eta_{t})}{\partial L^{(k)}_{t-1}},\kern 5.0ptk\in[K],\kern 5.0pt\text{with}\kern 5.0pt\mathcal{L}(\bm{L}_{t-1},\eta_{t}):=-\frac{1}{\eta_{t}}\log\left(\sum_{k=1}^{K}\exp(-\eta_{t}L^{(k)}_{t-1})\right)$,
where $\eta_{t}>0$ acts as a smoothing parameter. Large values of $\eta_{t}$ make the distribution over experts concentrate around the best expert(s) (recovering FTL), while small values yield nearly uniform weights. Equivalently, Hedge arises as the solution to an entropy-regularized linear optimization problem over the simplex, with Shannon entropy acting as a regularizer (see Boyd and Vandenberghe,, 2004). From this perspective, Hedge can be viewed as a smoothed version of FTL, with $\eta_{t}$ controlling the per-round smoothing of the loss.
Constant Hedge.
Setting $\eta_{t}:=\eta>0$ yields a constant-rate forecaster with a regret bound
$\overline{R}_{T}\leq\frac{\log K}{\eta}+\frac{\eta S_{T}^{2}T}{8},$ (7)
where $S_{T}$ is the maximal per-round loss range, which follows from Hoeffding’s lemma. Optimizing (7) with respect to $\eta$ gives
$\eta^{\star}=\sqrt{\frac{8\log K}{S_{T}^{2}T}},$
leading to $\overline{R}_{T}\leq S_{T}\sqrt{T\log K/2}$. If losses take values in $[0,1]$, then $S_{T}\leq 1$ and optimal $\eta^{\star}=\sqrt{8\log K/T}$ (Littlestone and Warmuth,, 1994). A constant rate allows for the standard closed-form online update via instantaneous losses:
$\omega^{(k)}_{\textnormal{{hdg}},t}=\begin{cases}1/K,&\text{if }t=1,\\[2.0pt] \dfrac{\omega^{(k)}_{\textnormal{{hdg}},t-1}\exp(-\eta\ell^{(k)}_{t-1})}{\sum_{j=1}^{K}\omega^{(j)}_{\textnormal{{hdg}},t-1}\exp(-\eta\ell^{(j)}_{t-1})},&\text{otherwise}.\end{cases}$
In practice, $\eta$ is a learning hyperparameter: $\eta\to\infty$ recovers FTL; $\eta=0$ yields no learning. The downside of a constant $\eta$ is its lack of adaptivity over arbitrarily long learning periods.
Decreasing Hedge (DecHedge).
Decreasing Hedge is obtained by setting $\eta_{t}=c_{0}\sqrt{\log K/t}$ for a constant $c_{0}>0$ (Auer and Chiang,, 2016). While this version of Hedge generally requires tuning $c_{0}$ for a specific learning run (as it is also the case for constant Hedge), Mourtada and Gaïffas, (2019) prove that the choice $c_{0}=2$ is worst-case optimal, without requiring further assumptions on the loss sequences. Decreasing Hedge, too, achieves a worst-case regret bound of $\mathcal{O}(\sqrt{T\log K})$ (Chernov and Zhdanov,, 2010). However, in stochastic settings where losses are well-behaved, more refined guarantees are possible. In fact, the generic constant $c_{0}=2$ may be conservative when the environment is not adversarial.
Adaptive Hedge (AdaHedge).
AdaHedge dynamically tunes $\eta_{t}$ based on past performance, and neither horizon nor additional parameters are used (van Erven et al.,, 2011).
Let $\bm{\omega}_{t}\in\Delta^{K-1}$ be the weight vector at time $t$, let $\bm{\ell}_{t}$ be the loss vector of $K$ experts, and let the forecaster’s loss be $\overline{\ell}_{t}$. The mix loss at time $t$ for a learning rate $\eta_{t}>0$ is
$\overline{m}_{t}:=-\frac{1}{\eta_{t}}\log\left(\bm{\omega}_{t}^{\top}\exp({-\eta_{t}\bm{\ell}_{t}})\right),\qquad\overline{M}_{t}:=\sum_{s=1}^{t}\overline{m}_{s},$
(see, e.g., Cesa-Bianchi and Lugosi,, 2006, Chapter 2). The approximation error $\delta_{t}:=\overline{\ell}_{t}-\overline{m}_{t}$ is the mixability gap, with cumulative gap $\nabla_{t}:=\sum_{s=1}^{t}\delta_{s}$. Consequently, the regret of the forecaster can be decomposed as
$\overline{R}_{t}=\overline{L}_{t}-\min_{k\in[K]}L^{(k)}_{t}=\overline{M}_{t}-\min_{k\in[K]}L^{(k)}_{t}+\nabla_{t},$
where $\overline{M}_{t}-\min_{k\in[K]}L^{(k)}_{t}$ is the regret incurred under the mix loss (de Rooij et al.,, 2014). Then, by (7),
$\overline{M}_{t}-\min_{k\in[K]}L^{(k)}_{t}+\nabla_{t}\leq\frac{\log K}{\eta}+\frac{\eta S_{t}^{2}t}{8}.$ (8)
AdaHedge chooses the learning rate at time $t$ by balancing the complexity term ${\log K}/{\eta}$ with the observed cumulative gap $\nabla_{t-1}$, [^3] At time $t$, while the forecaster makes their decision and the current loss is not realized yet, only the t1t-1 cumulative mixability gap is available. leading to:
$\eta_{\textnormal{{ah}},t}:=\frac{\log K}{\nabla_{t-1}},$
with the convention that $\eta_{\textnormal{{ah}},t}=\infty$ if $\nabla_{t-1}=0$. The scheme is thus parameter-free and independent of quantities that are available only at terminal time $T$.
Lemma 3.2 (de Rooij et al., 2014, Corollary 9).
The AdaHedge regret satisfies
$\overline{R}_{\textnormal{{ah}},T}\leq\sqrt{\sum_{t=1}^{T}s_{t}^{2}\log K}+S_{T}\left(\frac{4}{3}\log K+2\right).$
If $s_{t}\leq S$ for all $t$, this gives $\overline{R}_{\textnormal{{ah}},T}\leq S\sqrt{T\log K}+S\left(\frac{4}{3}\ln K+2\right)$, matching the $\mathcal{O}(S\sqrt{T\ln K})$ worst-case rate. Additionally, the leading term depends on the empirical variability $\sum_{t=1}^{T}s_{t}^{2}$ rather than directly on $T$, so the bound tightens in low-variance regimes.
Finally, for completeness, Appendix C also discusses the doubling-trick Hedge algorithm.
3.3 Online Learning Guarantees under Dependence
Following Luo and Schapire, (2015) and Mourtada and Gaïffas, (2019), one can see that the complexity of the online forecasting problem for a forecaster following a certain scheme can be generally understood in terms of the gaps between the expected loss of the best expert and the expected losses of other experts. Let the per-expert gap be defined as
$\Delta_{k}:=\mu_{k}-\mu_{k^{\star}}\geq 0,\kern 5.0ptk\in[K],$
where $\mu_{k}=\mathbb{E}[\ell^{(k)}_{t}$] and $k^{\star}\in\arg\min_{k\in[K]}\mu_{k}$, and let
$\Delta:=\min_{k\neq k^{\star}}\Delta_{k}=\min_{k\neq k^{\star}}\mu_{k}-\mu_{k^{\star}}$
denote the so-called sub-optimality gap between the expected loss of the best expert with respect to the runner-up.
3.3.1 Follow-the-Leader
Our first result is a set of regret guarantees for FTL online combination under both independence and mixing assumptions. We apply Hoeffding-type and Bernstein-type bounds to showcase the different kinds of guarantees that can be obtained, which depend on two constants that quantify the dependence of losses over time, and that are necessary ingredients in our proofs. For the sake of space, we do not include here an additional bound obtained using a Bernstein inequality which can be found in Appendix B.3. We emphasize that, even though the presented bounds assume a constant-in-time second moment for losses, they remain valid if one uses time-dependent variances or their uniform-in-time bound.
Theorem 3.1.
Let Assumptions 1-2 hold, and introduce $v_{\max}:=\max_{k\in[K]}v_{k}$, where $v_{k}:=\textnormal{Var}(\ell^{(k)}_{t})$, $k\in[K]$. Further assume that there are almost surely no ties in the FTL weights for all $t\geq 1$. Then, the expected regret of FTL is finite,
$\mathbb{E}\big[\overline{R}_{\textnormal{{ftl}},T}\big]\leq\min\big\{R^{H}_{\textnormal{{ftl}}},R^{B}_{\textnormal{{ftl}}}\big\}$
with the following cases:
If Assumption 2(i) holds, then
$R^{H}_{\textnormal{{ftl}}}=2+\frac{2\log K+4}{\Delta^{2}},\quad R^{B}_{\textnormal{{ftl}}}=2+\frac{(8v_{\max}+\tfrac{4}{3}\Delta)\left(\log(2K)+2\right)}{\Delta^{2}}.$
If Assumption 2(ii) and Assumption 4(i) hold, then
$R^{H}_{\textnormal{{ftl}}}$ $=3+\frac{8\overline{\theta}^{H}_{\max}(\log K+4)}{\Delta^{2}},\quad\text{with}\kern 5.0pt\overline{\theta}^{H}_{\max}=\max_{k\in[K]}\sup_{t\geq 1}\left(1+4\sum_{n=1}^{t-1}\varphi^{(k)}_{n}\right)$
$R^{B}_{\textnormal{{ftl}}}$ $=2+\frac{8\overline{\theta}^{B}_{\max}(8v_{\max}+\Delta)\left(\log(K)+2\right)}{\Delta^{2}}\quad\text{with}\kern 5.0pt\overline{\theta}^{B}_{\max}=\max_{k\in[K]}\sup_{t\geq 1}\left(1+\sum_{n=1}^{t}\sqrt{\varphi^{(k)}_{n}}\right)^{2}.$
Remark 3.1.
The $\Delta^{2}$ denominator in the bounds of Theorem 3.1 can, in fact, be sharpened in the i.i.d. setting (see Remark B.2 in the Appendix). This sharpening, however, hinges strictly on the independence of the loss process across time periods. An alternative approach that we do not pursue here would be to assume an upper bound on the time correlation weights and losses, as suggested by Gasparin and Ramdas, (2025).
3.3.2 Bounds for Hedge
Mourtada and Gaïffas, (2019) showed that, in the stochastic setting with i.i.d. losses, decreasing Hedge naturally adapts to the hardness of the combination problem, whereas constant and doubling trick Hedge do not.
Let $\chi^{(k)}_{t}:=(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})$ for each $\in[K], kk$$\not=k^{\star}$. The process $\{\chi^{(k)}_{t}\}_{t\in\mathbb{Z}}$ is the excess loss of expert $k$ relative to the best expert $k^{\star}$.
Theorem 3.2.
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss. Define $\widetilde{v}_{\max}:=\max_{k\neq k^{\star}}\textnormal{Var}(\chi^{(k)}_{t})$. Then, the expected regret of decreasing Hedge is finite,
$\mathbb{E}\big[\overline{R}_{\textnormal{{hdg}},T}\big]\leq\min\big\{R^{H}_{\textnormal{{hdg}}},R^{B}_{\textnormal{{hdg}}}\big\}$
with the following cases:
If Assumption 2(i) holds, then
$R^{H}_{\textnormal{{hdg}}}=\frac{4\Delta\log(K)+25}{\Delta^{2}},\quad R^{B}_{\textnormal{{hdg}}}=1+\sqrt{\log K}+\frac{4\sqrt{\frac{2}{3}}\Delta\log K+8(\widetilde{v}_{\max}+\frac{1}{3}\Delta)+16}{\Delta^{2}}.$
If Assumption 2(ii) and Assumption 3(ii) hold, then
$R^{H}_{\textnormal{{hdg}}}$ $=2+\frac{(1+3\overline{\rho}^{H}_{\max})(\Delta\log(K)+16)}{\Delta^{2}},$
with $\overline{\rho}^{H}_{\max}=\max_{k\neq k^{*}}\sup_{t\geq 1}\left(1+\sum_{n=1}^{t}\varphi_{n}(\{\chi^{(k)}_{t}\}_{t\in\mathbb{Z}})\right)$ and
$R^{B}_{\textnormal{{hdg}}}$ $=1+\sqrt{\log K}+\frac{4\sqrt{5}\overline{\rho}_{\max}^{B}\Delta\log K+16(\overline{\rho}_{\max}^{B})^{2}(4\widetilde{v}_{\max}+\Delta)+16}{\Delta^{2}},$
with $\overline{\rho}_{\max}^{B}=\max_{k\neq k^{\star}}\sup_{t\geq 1}\allowbreak\left(1+\sum_{n=1}^{t}\sqrt{\varphi_{n}\big(\{\chi^{(k)}_{t}\}_{t\in\mathbb{Z}}\big)}\right)$.
An additional but more complex bound can be constructed using a different Bernstein inequality, which, for clarity of exposition, we derive in Appendix B.3. Again, our Hedge regret bounds are valid for time-dependent variance or for time-independent bounds thereof.
3.4 Which Ensemble Method is Best for Forecasting?
As shown in Section 3.3, both Follow-the-Leader and decreasing Hedge algorithms admit constant regret upper bounds, for any terminal time $T$ and number of experts $K$. These guarantees are valid for an arbitrary choice of models-experts. This naturally leads to the question: What, in general, would be the best method for a time series forecasting setting?
The choice of a combination method is non-trivial. Theoretical upper bounds for cumulative regret depend on the sub-optimality gap $\Delta$, which depends on the unknown population mean of losses. Moreover, some methods, such as non-adaptive Hedge, typically require tuning of their learning rate, which could be problematic in practice. However, since in our non-adversarial stochastic setting where mixing dependence conditions are mildly different from independence, we can rely on numerical evidence from the literature to inform our selection. According to the numerical results in Mourtada and Gaïffas, (2019), FTL, decreasing Hedge, and AdaHedge all demonstrate strong performance, both for large $\Delta$-gap scenarios and “hard” stochastic zero-gap ($\Delta=0$) settings, considerably outperforming the constant and doubling-trick Hedge. However, the simulations by van Erven et al., (2011) show that constant and decreasing Hedge perform similarly in practice. In Section 5, we implement constant Hedge as a computationally simple baseline exponential combination method, decreasing Hedge with theoretically optimal scaling ($c_{0}=2$) and AdaHedge, which does not depend on any tuning parameter.
Remark 3.2.
In practice, one departs from the standard online learning setting in two relevant ways. First, the forecaster’s loss $\overline{\ell}_{t}:=\bm{\omega}_{t}^{\top}\bm{\ell}_{t}$ is a convex combination of expert losses, whereas the performance is often measured using the loss of the combined forecast, $\ell(\widehat{Y}_{t+h},Y_{t+h})$. If $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a convex map, then $\ell(\widehat{Y}_{t+h},Y_{t+h})\leq\overline{\ell}_{t}$, hence FTL and Hedge regret bounds remain informative. Second, we have adopted the standard assumption that losses are bounded, $\sup_{t\geq 1}\max_{k\in[K]}\lvert\ell^{(k)}_{t}\rvert\leq 1$. In applications, this requirement is typically mild, since, unless one works with heavy-tailed data, an appropriate normalization and a thresholding upper bound can be applied to unbounded losses, such as MSE or Huber loss.
4 Ensemble Echo State Networks
4.1 Echo State Networks in Brief
Echo State Networks (ESN) are a particular family of recurrent neural networks with randomly sampled connectivity weights. ESN models are, in general, nonlinear state-space systems that, in the forecasting setting, are defined by the following equations:
${X}_{t}$ $=\alpha X_{t-1}+(1-\alpha)\sigma(AX_{t-1}+CZ_{t}+{\zeta}),$ (9)
$Y_{t+1}$ $=b+W^{\top}X_{t}+\epsilon_{t+1},$ (10)
where $A\in\mathbb{R}^{D\times D}$ is the reservoir matrix, $C\in\mathbb{R}^{D\times d}$ is the input matrix, ${\zeta}\in\mathbb{R}^{D}$ is the input shift, $\alpha\in[0,1)$ is the leak rate, and ${\theta}:=(b,W^{\top})^{\top}\in\mathbb{R}^{D+1}$ with $W\in\mathbb{R}^{D}$ denotes the readout coefficients. The map $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activation function applied elementwise and we assume that $\sigma$ is the hyperbolic tangent in the rest of the paper. We refer to $A$, $C$, and $\zeta$} as state parameters. These parameters are randomly drawn and kept fixed throughout, while $\theta$ is the only set of parameters that need to be estimated.
The randomly drawn parameters $A$, $C$, and $\zeta$} in (9) are constructed by first sampling $\widetilde{A}$, $\widetilde{C}$, and $\widetilde{{\zeta}}$ from appropriately chosen distributions, and then normalizing as follows:
$\overline{A}={\widetilde{A}}/{\rho(\widetilde{A})},\quad\overline{C}={\widetilde{C}}/{\|\widetilde{C}\|},\quad\overline{{\zeta}}={\widetilde{{\zeta}}}/{\|\widetilde{{\zeta}}\|}$, (11)
[…]
$X_{t}=\alpha X_{t-1}+(1-\alpha)\sigma(\rho\overline{A}X_{t-1}+\gamma\overline{C}Z_{t}+\varsigma\overline{{\zeta}}).$ (12)
The tuple ${\varphi}:=(\alpha,\rho,\gamma,\varsigma)$ is referred to as the hyperparameters of the ESN. Specifically, $\alpha\in[0,1)$ is the so-called leak rate, $\rho\in\mathbb{R}^{+}$ denotes the spectral radius of the reservoir matrix $A$, $\gamma\in\mathbb{R}^{+}$ is the input scaling, and $\varsigma\in\mathbb{R}^{+}$ corresponds to the shift scaling. These hyperparameters play a crucial role in defining the properties of the state map in (12). Although there is no clear consensus on how to best select $\varphi$, in this work we construct ensembles of ESN models that not only differ in the draws of the state parameters $A$, $C$, and $\zeta$, but also in the state hyperparameters. This allows us to explore a wide range of different reservoir architectures, and thus sidestep the problem of model tuning. [^4] Ballarin et al., 2024a propose a general data-driven algorithm to tune $\varphi$. However, the resulting hyperparameter optimization problem is computationally complex and highly non-convex.
As is common in the ESN literature, we consider the ridge regression estimator for $\theta$. More precisely, let $X:=(X_{1},\ldots,X_{T-1})^{\top}\in\mathbb{R}^{(T-1)\times D}$ and $Y:=(Y_{2},\ldots,Y_{T})^{\top}\in\mathbb{R}^{T-1}$. Define $M=\big(\mathbb{I}_{T-1}-\frac{1}{T-1}\mathbf{1}_{T-1}\mathbf{1}_{T-1}^{\top}\big)$, $\widetilde{Y}=MY$, and $\widetilde{X}=MX.$ Then the ridge regression estimator for $\theta$ is given by
$\widehat{W}_{\lambda}$ $=\arg\min_{W\in\mathbb{R}^{D}}\{\left\|\widetilde{Y}-\widetilde{X}W\right\|_{2}^{2}+\lambda\left\|W\right\|_{2}^{2}\}=\big(\widetilde{X}^{\top}\widetilde{X}+\lambda\mathbb{I}_{D}\big)^{-1}\widetilde{X}^{\top}\widetilde{Y}$, (13)
$\widehat{b}_{\lambda}$ $=\frac{1}{T-1}{\bf{1}}_{T-1}^{\top}\left(Y-X\widehat{W}_{\lambda}\right),$ (14)
where $\lambda\in\mathbb{R}^{+}$ is the regularization hyperparameter. In practice, to select $\lambda$ we rely on time-series-adapted cross-validation (CV) (see, e.g., Bergmeir and Benitez,, 2012; Bergmeir et al.,, 2018; Hyndman and Athanasopoulos,, 2018).
4.2 Multi-Frequency ESNs
Ballarin et al., 2024a introduced a class of Echo State Network models designed for time series observed at multiple sampling frequencies. These models are collectively referred to as Multi-Frequency ESNs (MFESNs).
The first category of MFESNs, known as Single-reservoir MFESNs (S-MFESNs), enables the forecasting of lower-frequency targets by mapping high-frequency reservoir states to low-frequency outputs through a state alignment scheme. In an aligned S-MFESN, the forecast is generated based on the most recent state corresponding to the reference low-frequency time index $t$. The second category, termed Multi-reservoir MFESNs (M-MFESNs), introduces multiple state equations, each associated with a distinct sampling frequency in the input data. For example, quarterly and monthly variables can be incorporated jointly as regressors. By employing multiple reservoirs evolving at their respective frequencies, these models flexibly capture the dynamics of different groups of input variables sampled at common frequencies.
We now briefly introduce the M-MFESN framework following Ballarin et al., 2024a . For brevity, we do not discuss S-MFESN models explicitly, as they are empirically dominated by multi-reservoir MFESNs (see Section 5). Suppose there are $Q$ groups of input time series with observations $\{Z^{(q)}_{t,{s}|{\kappa_{q}}}\}_{t,s}$, where $Z^{(q)}_{t,{s}|{\kappa_{q}}}\in\mathbb{R}^{d_{q}}$, $\kappa_{q}\in\mathbb{N}$, $q\in[Q]$, $t\in[T]$, and $s\in\{0,\ldots,\kappa_{q}-1\}$, sampled at common frequencies $\{\kappa_{1},\ldots,\kappa_{Q}\}$. Using the tempo indexing of Ballarin et al., 2024a , the high-frequency observations occur at fixed fractional sub-intervals relative to the reference low-frequency index $t$. For each frequency $\kappa_{q}$, the fractional index $s=0$ indicates alignment with $t$, so that $(t,{0}|{\kappa_{q}})\equiv t$, and advancing $\kappa_{q}$ sub-intervals corresponds to one unit of low-frequency time, that is $(t,{\kappa_{q}}|{\kappa_{q}})\equiv(t+1,{0}|{\kappa_{q}})\equiv t+1$. For each input series group $q$, we define a frequency-specific $D_{q}$-dimensional reservoir state equation
$X^{(q)}_{t,{s}|{\kappa_{q}}}=\alpha_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+(1-\alpha_{q})\sigma(\rho_{q}\overline{A}_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+\gamma_{q}\overline{C}_{q}Z^{(q)}_{t,{s}|{\kappa_{q}}}+\varsigma_{q}\overline{\zeta}_{q}),$ (15)
where $A_{q}$, $C_{q}$, and $\zeta_{q}$ are group-specific normalized reservoir parameters, and ${\varphi}_{q}:=(\alpha_{q},\rho_{q},\gamma_{q},\varsigma_{q})$ is the corresponding hyperparameter vector of the reservoir. To construct the forecast at the reference time $t$, the states are aligned by stacking the frequency-specific states evaluated at $s=0$, that is, $X_{t,Q}:=(X^{(1)\top}_{t,{0}|{\kappa_{1}}},\ldots,X^{(Q)\top}_{t,{0}|{\kappa_{Q}}})^{\top}\in\mathbb{R}^{D_{Q}}$, $D_{Q}:=\sum_{q=1}^{Q}D_{q}$. The M-MFESN prediction equation is given by
$Y_{t+1}$ $=b_{Q}+W_{Q}^{\top}X_{t,Q}+\epsilon_{t+1}=b_{Q}+\sum_{q=1}^{Q}W_{q}^{\top}X^{(q)}_{t,{0}|{\kappa_{q}}}+\epsilon_{t+1}.$ (16)
The readout $W_{Q}$ defines a linear map from the nonlinear, frequency-specific states to $Y_{t+1}$.
4.3 Ensemble (Multi-Frequency) ESNs
An important concern when working with models initialized or inherently constructed with random weights is the impact that such randomness may have on model performance. A recent and growing body of work on the “lottery ticket” hypothesis (Frankle and Carbin,, 2018; Malach et al.,, 2020), argues, informally, that a key component of the empirical success of complex neural networks is “being lucky”, that is, drawing an initial configuration of weights that is favorable (Ma et al.,, 2021; Sreenivasan et al.,, 2022). Related research has focused explicitly on designing weight distributions that are guaranteed to be a priori advantageous for initialization (see, e.g., Zhao et al., 2022 and Bolager et al., 2023). Echo State Networks, although proven to satisfy universal approximation and generalization guarantees (Grigoryeva and Ortega, 2018a, ; Gonon et al., 2020b, ; Gonon et al.,, 2023), also depend critically on the properties of randomly drawn reservoir parameters. Once used for forecasting such systems may offer very different forecasting accuracy. It is hence suggestive to construct ensembles of ESNs (EN-ESN). In our multifrequency setting we directly extend this idea to ensembles of MFESNs (EN-MFESN), with single- or multi-reservoir architectures.
Following the notation introduced in Section 4.2, consider an ensemble of $K\geq 1$ distinct MFESN models-experts of multiple reservoirs (M-MFESNs), each used to construct the corresponding expert prediction out of $Q$ groups of input series sampled at distinct frequencies as $\widehat{Y}^{(k)}_{t+1}=\widehat{b}^{(k)}_{Q,\lambda}+\widehat{W}^{(k)\top}_{Q}X^{(k)}_{t,Q}$, $k\in[K]$, with $\widehat{W}^{(k)\top}_{Q}$ and $\widehat{b}^{(k)}_{Q,\lambda}$ estimated as in (13)-(14). The ensemble forecast is constructed according to (2) as
$\widehat{Y}_{t+1}=\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{Y}^{(k)}_{t+1}=\widehat{b}_{K}+\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{W}^{(k)\top}_{Q}X^{(k)}_{t,Q},\kern 5.0pt{\rm with}\kern 5.0pt\widehat{b}_{K}:=\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{b}^{(k)}_{Q,\lambda}$. (17)
ESN Ensembles over Random Parameters (EN-MFESN-RP).
Since, as discussed above, the primary source of diversity MFESN instances is the randomness of the state parameters, the first ensemble class we consider is obtained by re-drawing the state parameters for each $k\in[K]$, as in (15), while keeping all other reservoir hyperparameters $\varphi_{q}=(\alpha_{q},\rho_{q},\gamma_{q},\varsigma_{q})$ fixed (for single- and multi-reservoir cases). This type of reservoir ensemble, which we call MFESN-RP ensemble (EN-MFESN-RP), aims to span the family of models that differ only in terms of the sampled state parameters. As a result, only moderate variation can be expected between members in terms of autoregressive state dynamics, input scaling, and possible activation function saturation.
[…]
5 GDP Forecasting with EN-MFESNs
In this section, we evaluate and compare the one-step-ahead forecasting performance of our proposed EN-MFESN against individual MFESN models and state-of-the-art approaches. To ensure a fair comparison, we use the same empirical experimental setup as in Ballarin et al., 2024a : We use the same data, preprocessing procedures, and benchmark specifications.
5.1 Data
We conduct our analysis for the original dataset used in Ballarin et al., 2024a , which contains two groups of predictors: small-MD, comprising 9 predictors, and medium-MD, consisting of 33 predictors. The data is collected at different sampling frequencies, while the target variable is U.S. quarterly GDP growth. In this paper, we focus on a more challenging medium-MD data sample, while the results for the small-MD dataset can be made available by request. For more in-depth information on specific data sources, data preprocessing and availability, we refer the reader to Appendix F and Section 4.1 in Ballarin et al., 2024a .
5.2 Models and Ensembles
Benchmarks.
The first benchmark we consider is the in-sample mean of U.S. GDP growth computed over the estimation sample, which serves a natural baseline for comparing mean squared forecasting errors. We also include an AR(1) model, standard in univariate time series analysis and macroeconomic forecasting (Stock and Watson,, 2002; Bai and Ng,, 2008). Our analysis demonstrates that AR(1) can outperform naïve forecast and poorly calibrated ensembles, making it a non-trivial reference for evaluating multi-frequency nonlinear state-space models. Additionally, the two dynamic factor models (DFMs) from Ballarin et al., 2024a are also included. DFMs are widely applied in macroeconomic analysis to recover low-dimensional common components from high-dimensional time series (Stock and Watson,, 2016; Doz and Fuleky,, 2020; Barigozzi and Hallin,, 2024).
MFESNs.
Two classes of ESN-based forecasting models are considered. The first set comprises single-frequency ESN models (S-MFESNs) with reservoir sizes $D=30$ and $D=120$, respectively. The sparsity degree of state matrices is set to $10/D$. The second group of models includes multi-frequency ESNs (M-MFESNs) with separate reservoirs for monthly and daily data input series, allowing to capture the dynamics across both temporal resolutions. For each model class, two variants of MFESNs are considered: “A” and “B” types, which differ, e.g., in the state hyperparameters used. Our implementation of MFESNs follows Ballarin et al., 2024a in the choice of architecture, reservoir sizes, activation functions, hyperparameters, time-series cross validation for the regularization parameter, and other relevant model-specific details. A summary of our model designs is provided in Appendix I.
Ensembles.
[…]
(a) S-MFESN A
(b) S-MFESN B
(c) M-MFESN A
(d) M-MFESN B
Figure 1: EN-MFESN-RP ensembles constructed out of S-MFESN A (a), S-MFESN B (b), M-MFESN A (c), and M-MFESN B (d) models. Plots display MSFE relative to AR(1) baseline for ensemble aggregation methods (colored dashed lines), the median ensemble performance (black dotted line), specific draws of S-MFESN A/B and M-MFESN A/B from Ballarin et al., 2024a (black solid line), and empirical CDF of the relative MSFE across individual ensemble models (gray curve). Colored markers are included to distinguish overlapping lines.
(a) S-MFESN A
(b) S-MFESN B
(c) M-MFESN A
(d) M-MFESN B
[…]
5.3 Forecasting Performance
[…]
Figure 1 presents the results for EN-MFESN-RP ensembles and reveals substantial sensitivity to the random draw of reservoir state coefficients. In particular, individual S-MFESN A/B and M-MFESN A/B model draws taken from Ballarin et al., 2024a display heterogeneous performance: While B-type single- and multi-reservoir MFESNs perform favorably, A-type models often exhibit higher forecasting errors, in some cases exceeding the median relative MSE with respect to the AR(1) benchmark. Simple averaging and constant Hedge follow each other closely across panels, primarily due to the choice of a small learning rate. In contrast, both FTL and AdaHedge deliver consistent, sizable, and robust performance gains. It is important to notice that Follow-the-Leader often to models in the extreme left tail of the ECDF, indicating near-optimal performance within the ensemble. Although FTL dominates AdaHedge throughout, the difference is higher for S-MFESN models than for M-MFESNs, consistent with their greater flexibility and baseline accuracy of the latter.
[…]
Ensemble
Model Baseline Median Average RollMSE FTL Hedge DecHedge AdaHedge
Mean 1.000
AR(1) 0.758
DFM A 0.835
DFM B 1.093
EN-MFESN-RP (random parameters resampling)
S-MFESN A 0.970 0.940 0.921 0.917 0.661 0.844 0.838 0.714
-3.04% -5.07% -5.40% -31.85% -13.00% -13.54% -26.34%
S-MFESN B 0.832 0.888 0.880 0.878 0.616 0.831 0.812 0.676
+6.68% +5.74% +5.44% -26.01% -0.21% -2.40% -18.76%
M-MFESN A 0.903 0.897 0.891 0.890 0.736 0.872 0.870 0.756
-0.58% -1.28% -1.42% -18.43% -3.41% -3.57% -16.20%
M-MFESN B 0.683 0.737 0.731 0.727 0.588 0.689 0.682 0.599
+7.96% +7.00% +6.49% -13.95% +0.91% -0.13% -12.35%
[…]
S-MFESN A 0.970 0.897 0.839 0.822 0.676 0.665 0.631 0.553
-7.46% -13.50% -15.23% -30.34% -31.42% -34.94% -42.95%
S-MFESN B 0.832 0.837 0.783 0.767 0.488 0.640 0.617 0.507
+0.56% -5.91% -7.84% -41.41% -23.10% -25.83% -39.08%
M-MFESN A 0.903 0.902 0.876 0.868 0.634 0.742 0.707 0.601
-0.10% -3.01% -3.90% -29.81% -17.84% -21.71% -33.44%
M-MFESN B 0.683 0.794 0.761 0.745 0.481 0.647 0.626 0.529
+16.32% +11.42% +9.02% -29.60% -5.20% -8.37% -22.47%
Table 2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing combinations are displayed in bold.
[…]
5.4 Hyperparameter Sensitivity
[…]
(a) S-MFESN A
(b) S-MFESN B
(c) M-MFESN A
(d) M-MFESN B
[…]
These findings suggest revisiting the leak-rate choices in Ballarin et al., 2024a . Across MFESN classes, the most effective leak rates differ substantially from the original specifications. For instance, baseline S-MFESNs are constructed with $\alpha=0.1$ (Table I.1), while Figure 3(a)-(b) suggest the values around $0.7$ are a uniformly more robust choice. This insight, however, comes at a cost. Attempting to study multiple hyperparameters with a comparable level of precision, such as distinct monthly and daily frequency reservoir leak rates in M-MFESNs, would require the ensemble budget to grow exponentially. Model combination therefore offers a parsimonious and adaptive alternative, relying on a fixed-budget set of models rather than the construction of a single optimal specification. This adaptivity is particularly valuable for MFESNs, where randomness induces persistent model heterogeneity that cannot be corrected ex-post, a feature shared by many modern neural network architectures.
6 Conclusion
The ensemble combination approach is a well-studied framework that enables a forecaster to incorporate the predictions of a pool of models into a single, adaptive forecast. However, these techniques remain under-explored in the fields of applied statistics and econometrics.
In this setting, our work makes contributions in both theoretical and empirical directions. First, we provide concentration bounds for stochastic ensemble losses under i.i.d. and $\varphi$-mixing assumptions, exploiting Hoeffding- and Bernstein-type results. Then, we apply these results to Follow-the-Leader and decreasing Hedge methods, for which we derive finite regret bounds. We then consider an application where model combination allows for robust and substantial performance improvements against previous state-of-the-art results: Macroeconomic forecasting with mixed-frequency data using multi-frequency ESN models. We find that the ensemble approach is extremely effective at sharpening the performance of all MFESN models, with Follow-the-Leader and AdaHedge schemes being the overall most effective schemes. Our empirical analysis also reveals that combinations can efficiently sidestep the hyperparameter tuning, which for (MF)ESN models is highly non-trivial.
Appendix
From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles
Giovanni Ballarin     Lyudmila Grigoryeva     Yui Ching Li
Appendix A Technical Lemmas
A.1 Basic Results
Proposition A.1 (Chernov and Zhdanov,, 2010).
Let $\eta_{1},\eta_{2},\ldots$ be a decreasing sequence of learning rates. The Hedge algorithm satisfies the following regret bound:
$\overline{R}_{\textnormal{{hdg}},T}\leq\frac{1}{\eta_{T}}\log K+\frac{1}{8}\sum_{t=1}^{T}\eta_{t}.$ (18)
The choice of $\eta_{t}=2\sqrt{\log K/t}$ yields a regret bound of $\sqrt{T\log K}$ for every $T\geq 1$.
Lemma A.1.
Let $a,b>0$ be constants. Then for any $x\geq 0$
$e^{-x/a}+e^{-x/b}\leq 2e^{-{x}/{\max\{a,b\}}}.$
Lemma A.2 (Mourtada and Gaïffas,, 2019, Lemma 13).
For every $\alpha>0$,
$\sum_{t=1}^{\infty}e^{-\alpha t}\leq\frac{1}{\alpha}\quad\text{and}\quad\sum_{t=1}^{\infty}e^{-\alpha\sqrt{t}}\leq\frac{2}{\alpha^{2}}.$
Lemma A.3.
Let $b>0$ and $\alpha>0$ be constants. Then, there exists $\nu^{*}>0$ such that, setting $t^{*}=\nu^{*}b$, it holds
$\frac{t^{*}}{(\log t^{*})^{\alpha}}\geq b.$
Consider the function $g(t)=t/(\log(t))^{\alpha}$, $t>1$. For all $t\geq e^{\alpha}$ is strictly increasing and continuous. Now, letting $t=\nu b$ for some $\nu\geq b$, it clearly holds
$g(t)=\frac{\nu b}{(\log\nu+\log b)^{\alpha}}\geq\frac{\nu b}{(2\log\nu)^{\alpha}}$ (19)
since the denominator of the left-hand side is smaller than on the right-hand side by the lower bound assumed on $\nu$. Hence, we can define
$\nu^{*}:=\inf\left\{\nu\geq\max\{e^{\alpha},b\}\,\big|\,\nu\geq(2\log\nu)^{\alpha}\right\}.$
By continuity of $g(t)$, $\nu^{*}$ is also in the set, that is,
$\frac{\nu^{*}}{(2\log\nu^{*})^{\alpha}}\geq 1$
and hence
$\frac{\nu^{*}b}{(2\log\nu^{*})^{\alpha}}\geq b.$
Setting $t^{*}=\nu^{*}b$, and using (19), we obtain $g(t^{*})\geq b$ as required. ∎
A.2 Weak Dependence
First, we recall the formal definition of $\varphi$-mixing, following (Rio,, 2017).
Definition A.1 (Uniform mixing coefficients, Rio,, 2017, Definition 2.1).
Let $\{U_{i}\}_{i\in\mathbb{Z}}$ be a sequence of real-valued random variables. The uniform mixing coefficients of $\{U_{i}\}_{i\in\mathbb{Z}}$ are defined by
$\varphi_{0}:=1\quad\text{and}\quad\varphi_{n}:=\sup_{m\in\mathbb{Z}}\varphi(\mathcal{F}_{m},\sigma(U_{m+n}))\text{ for any }n\in\mathbb{N}.$
The sequence $\{U_{i}\}_{i\in\mathbb{Z}}$ is said to be uniformly mixing if $\lim_{n\to\infty}\varphi_{n}=0$.
Hoeffding-type bounds.
The book of Rio, (2017) provides a generalization of the classical i.i.d. version of Hoeffding’s inequality to $\varphi$-mixing processes.
Lemma A.4 (Hoeffding inequality for uniformly mixing, Rio,, 2017, Corollary 2.1).
Let $\{U_{i}\}_{i\geq 0}$ be a sequence of centered and real-valued bounded random variables such that $\lvert U_{i}\rvert^{2}\leq M_{i}$. Set $\theta_{n}:=1+4\sum_{i=1}^{n-1}\varphi_{i}$. Then, for any positive $\xi$,
$\mathbb{P}\left(\left\lvert\sum_{i=1}^{n}U_{i}\right\rvert\geq\xi\right)\leq\sqrt{e}\exp\left(-\frac{\xi^{2}}{2\theta_{n}\sum_{i=1}^{n}M_{n}}\right).$
Bernstein-type bounds.
Bernstein’s inequality provides a more general bound than Hoeffding’s by taking into account the second moments of the random variables involved.
Lemma A.5 (Bernstein inequality for i.i.d., Boucheron et al.,, 2013, Corollary 2.11).
Let $\{U_{i}\}_{i\geq 0}$ be a sequence of centered and real-valued bounded random variables such that $\lvert U_{i}\rvert\leq M$ for all $i$. Let $\textnormal{Var}(U_{i})=v<\infty$ for all $i$. Then for any $\xi>0$,
$\mathbb{P}\left(\sum_{i=1}^{n}U_{i}\geq\xi\right)\leq\exp\left(-\frac{\xi^{2}}{2nv+\frac{2}{3}M\xi}\right).$
We recall here two generalizations of classical Bernstein which apply to the $\varphi$-mixing setting, due respectively to Samson, (2000) and Hang, (2015); Hang and Steinwart, (2017).
Lemma A.6 (Bernstein inequality for $\varphi$-mixing).
Let $\{U_{i}\}_{i\geq 0}$ be a real-valued stationary process such that $\lvert U_{i}\rvert\leq M$ for all $i$. Let $\varphi_{i}$ denote its $\varphi$-mixing coefficients and let $\textnormal{Var}(U_{i})=v<\infty$ for all $i$. Define $\theta^{2}_{n}:=\big(1+\sum_{i=1}^{n}\sqrt{\varphi_{i}}\big)^{2}$ and
$Z:=\left\lvert\,\sum_{i=1}^{n}U_{i}\,\right\rvert$
Then for any $\xi>0$,
$\mathbb{P}\left(Z\geq\mathbb{E}[Z]+\xi\right)\leq\exp\left(-\frac{1}{8\theta^{2}_{n}}\min\left\{\frac{\xi}{M},\ \frac{\xi^{2}}{4nv}\right\}\right)\leq\exp\left(-\frac{\xi^{2}}{4\theta^{2}_{n}(4nv+M\xi)}\right)$ (20)
and
$\mathbb{P}\left(Z\leq\mathbb{E}[Z]-\xi\right)\leq\exp\left(-\frac{1}{8\theta^{2}_{n}}\min\left\{\frac{\xi}{M},\ \frac{\xi^{2}}{4nv}\right\}\right)\leq\exp\left(-\frac{\xi^{2}}{4\theta^{2}_{n}(4nv+M\xi)}\right).$ (21)
The first inequalities in (20) and (21) follow trivially from Samson, (2000), Theorem 3, where the function applied to the random variables, $g$, is the identity. Then, $\lvert U_{i}\rvert\leq M$ trivially ensures the boundedness of $g(U_{i})$ and
$\mathbb{E}\left[\sum_{i=1}^{n}g(U_{i})^{2}\right]\leq nv,$
according to the assumptions in p. 453 of Samson, (2000). The calculation of factor $\theta^{2}_{n}$, which is related to the operator norm of the dependence matrix $\Gamma$ in the theorem, also follows from the discussion in Samson, (2000), pp. 420-425 (see also Alquier et al.,, 2013). The second inequalities in (20) and (21) directly follow by noticing that for any $a,b\in\mathbb{R}$ it holds that $\min\{a,b\}\geq H(a,b)$ with H(a,b)=2ab/(a+b)H(a,b)={2ab}/{(a+b)} the harmonic mean of $a$ and $b$, and hence
$\min\left\{\frac{\xi}{M},\ \frac{\xi^{2}}{4nv}\right\}\geq\frac{2\xi^{2}}{4nv+M\xi}.$
The next result we present holds for the wider class of $\mathcal{C}$-mixing processes, and can be immediately adapted to $\varphi$-mixing sequences (see Hang,, 2015, Chapter 4 for a detailed discussion). The price to pay for this generalization is, however, an additional logarithmic factor in the denominator of the exponential which is tied to the dependence structure of the process.
Theorem A.1 (Bernstein inequality for geometrically $\varphi$-mixing, Hang,, 2015, Theorem 4.7).
Let $\{U_{i}\}_{i\geq 0}$ be a stationary, geometrically $\varphi$-mixing process with rate $(d_{n})_{n\geq 0}$ of the form $d_{n}\leq c\exp{(-bn^{\gamma})}$, $n\geq 1$, for some $b>0$, $c\geq 0$, and $\gamma>0$. Let $h:\mathbb{R}\to\mathbb{R}$ be bounded, measurable with $\|h\|_{\infty}\leq M$, $\mathbb{E}[h]=0$, and $\textnormal{Var}(h)\leq\sigma^{2}$. Then there exists
$n_{0}=\max\big\{\min\left\{m\geq 3\>|\>m^{2}\geq{3232cM}\textnormal{ and }{m}{(\log m)^{-2/\gamma}}\geq 4\right\},e^{3/b}\big\}$
such that, for all $n\geq n_{0}$ and all $\varepsilon>0$,
$\mathbb{P}\left(\sum_{i=1}^{n}h(U_{i})\geq\varepsilon\right)\leq 2\exp\left(-\frac{\varepsilon^{2}}{8(\log n)^{2/\gamma}\left(\sigma^{2}+\varepsilon M/3\right)}\right).$
Appendix B Proofs
This section contains the proofs for the theoretical results in the paper, as well as all intermediate technical results needed to derive them. First, we present the main concentration results for the experts’ losses; then, we obtain regret bounds for FTL and Hedge methods.
B.1 Concentration Results for FTL
We exploit the theory from Appendix A.2 to construct concentration-type bounds for the cumulative losses of experts under both i.i.d. and dependent settings.
B.1.1 Hoeffding-type
Lemma B.1 (Hoeffding-type).
Let Assumptions 1-2 hold and, without loss of generality, assume that the experts are ordered such that $\mu_{1}<\mu_{2}\leq\ldots\leq\mu_{K}$, that is $\Delta=\mu_{2}-\mu_{1}>0$.
If Assumption 2(i) holds, then
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq\exp\left(-\frac{t\Delta^{2}}{2}\right)\left[2+\sum_{k=3}^{K}\exp\left(-2t(\mu_{1}-\mu_{k})(\mu_{2}-\mu_{k})\right)\right].$
If Assumption 2(ii) holds, then
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq\sqrt{e}\sum_{k=1}^{K}\exp\left(-\frac{t(\mu_{1}-\mu_{k}+\Delta/2)^{2}}{2\theta^{(k)}_{t}}\right)$,
where $\theta^{(k)}_{t}:=1+4\sum_{n=1}^{t-1}\varphi^{(k)}_{n}$.
First, choose some $\delta$ such that $\mu_{1}<\delta<\mu_{2}$. For notational convenience, define $L^{*\setminus 1}_{t}:=\min_{2\leq k\leq K}L^{(k)}_{t}$ and event $\mathcal{A}:=\{L^{*\setminus 1}_{t}\leq t\delta\}$, whose complement is denoted by $\mathcal{A}^{\complement}$.
By the law of total probability,
$\mathbb{P}(L^{(1)}_{t}>L^{*\setminus 1}_{t})$ $=\mathbb{P}\left(\{L^{(1)}_{t}>L^{*\setminus 1}_{t}\}\cap\mathcal{A}\right)+\mathbb{P}\left(\{L^{(1)}_{t}>L^{*\setminus 1}_{t}\}\cap\mathcal{A}^{\complement}\right)$
$\leq\mathbb{P}(\mathcal{A})+\mathbb{P}\left(\{L^{(1)}_{t}>L^{*\setminus 1}_{t}\}\cap\mathcal{A}^{\complement}\right).$
To upper bound the first term, since $K$ is finite and fixed, one can directly apply the union bound and obtain
$\mathbb{P}\left(\min_{2\leq k\leq K}L^{(k)}_{t}\leq t\delta\right)$ $\leq\sum_{k=2}^{K}\mathbb{P}\left(L^{(k)}_{t}\leq t\delta\right)=\sum_{k=2}^{K}\mathbb{P}\left(L^{(k)}_{t}-t\mu_{k}\leq t(\delta-\mu_{k})\right)$
$=\sum_{k=2}^{K}\mathbb{P}\left(\sum_{s=1}^{t}(\ell^{(k)}_{s}-\mu_{k})\leq t(\delta-\mu_{k})\right).$
Case (i). We first consider the case where Assumption 2(i) holds. By Hoeffding’s inequality, since $\ell^{(k)}_{t}-\mu_{k}\in[-\mu_{k},1-\mu_{k}]$, for $k\neq 1$ it holds
$\mathbb{P}\left(\sum_{s=1}^{t}(\ell^{(k)}_{s}-\mu_{k})\leq tm_{k}\right)$ $\leq\exp\left(-2tm_{k}^{2}\right),$
where $m_{k}=\delta-\mu_{k}$. Hence,
$\mathbb{P}(\mathcal{A})\leq\sum_{k=2}^{K}\exp\left(-2t(\delta-\mu_{k})^{2}\right).$
Since
$\mathbb{P}\left(\{L^{(1)}_{t}>L^{*\setminus 1}_{t}\}\cap\{L^{*\setminus 1}_{t}>t\delta\}\right)\leq\mathbb{P}\left(L^{(1)}_{t}>t\delta\right),$
by applying Hoeffding’s inequality again we obtain
$\mathbb{P}\left(L^{(1)}_{t}>t\delta\right)\leq\exp\left(-2t(\delta-\mu_{1})^{2}\right).$
Set now $\delta=(\mu_{1}+\mu_{2})/2=\mu_{1}+\Delta/2$, so that
$\mathbb{P}\left(L^{(1)}_{t}>L^{*\setminus 1}_{t}\right)$ $\leq 2\exp\left(-t\Delta^{2}/2\right)+\sum_{k=3}^{K}\exp\left(-2t(\mu_{1}-\mu_{k}+\Delta/2)^{2}\right)$
$=\exp\left(-t\Delta^{2}/2\right)\left[2+\sum_{k=3}^{K}\exp\left(-2t(\mu_{1}-\mu_{k})(\mu_{1}-\mu_{k}+\Delta)\right)\right].$
The final bound follows by using $\Delta=\mu_{2}-\mu_{1}$.
Case (ii). When working under mixing conditions, we can leverage known generalizations of Hoeffding’s inequality to dependent data. Under Assumption 2(ii), the Hoeffding-type bound of Lemma A.4 (see also Rio,, 2017) in our setting states that, for any $\xi>0$,
$\mathbb{P}\left(\left\lvert\,L^{(k)}_{t}-\mu_{k}\,\right\rvert\geq\xi\right)\leq\sqrt{e}\exp\left(-\frac{\xi^{2}}{2\theta^{(k)}_{t}t}\right).$
We follow the same arguments as in the above proof for case (i), and find that
$\mathbb{P}\left(\sum_{s=1}^{t}(\ell^{(k)}_{s}-\mu_{k})\leq tm_{k}\right)$ $\leq\sqrt{e}\exp\left(-\frac{tm_{k}^{2}}{2\theta^{(k)}_{t}}\right)$
and
$\mathbb{P}\left(L^{(1)}_{t}>t\delta\right)\leq\sqrt{e}\exp\left(-\frac{t(\delta-\mu_{1})^{2}}{2\theta^{(1)}_{t}}\right)$.
Combining these bounds and choosing once more $\delta=\mu_{1}+\Delta/2$ concludes the proof. ∎
Corollary B.1.
Let the conditions of Lemma B.1 hold. Then the following bounds hold:
Under the conditions of case (i) of Lemma B.1
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq K\exp\left(-t\frac{\Delta^{2}}{2}\right).$ (22)
Under the conditions of case (ii) of Lemma B.1, assume, additionally, that Assumption 3(i) holds and define $\overline{\theta}_{\max}=\max_{k\in[K]}\sup_{t\geq 1}\theta^{(k)}_{t}$. Then
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq K\sqrt{e}\exp\left(-\frac{t\Delta^{2}}{8\overline{\theta}_{\max}}\right).$ (23)
Part (i) is an immediate consequence of case (i) of Lemma B.1 using that, since $\left(\mu_{1}-\mu_{k}\right)\left(\mu_{2}-\mu_{k}\right)\geq 0$ for all $k\geq 3$, then $\exp\left(-2t\left(\mu_{1}-\mu_{k}\right)\left(\mu_{2}-\mu_{k}\right)\right)\leq 1$ for all $k\geq 3$, and hence $\left[2+\sum_{k=3}^{K}\exp\left(-2t(\mu_{1}-\mu_{k})(\mu_{2}-\mu_{k})\right)\right]<K$. Part (ii) is obtained directly from case (ii) of Lemma B.1 by noticing that $\Delta/2\leq\mu_{1}-\mu_{k}+\Delta/2\leq\Delta/2$ and $\theta_{t}^{(k)}\leq\bar{\theta}_{\max}$, for all $k\in[K]$ and for all $t$, where by Assumption 3(i), $\bar{\theta}_{\max}\leq 1+4C_{1,\varphi}<\infty$. ∎
B.1.2 Bernstein-type
Lemma B.2 (Bernstein-type).
Let Assumptions 1-2 hold and let $v_{k}:=\textnormal{Var}(\ell^{(k)}_{t})$, $k\in[K]$.
If Assumption 2(i) holds, then for all $t\geq 1$,
$\mathbb{P}\left(L^{(k^{\star})}_{t}>\min_{k\neq k^{\star}}L^{(k)}_{t}\right)$ $\leq 2\sum_{k\neq k^{\star}}\exp\left(-\frac{t\Delta_{k}^{2}}{8\max\{v_{k},v_{k^{\star}}\}+\tfrac{4}{3}\Delta_{k}}\right).$ (24)
If Assumption 2(ii) holds, define ${\theta}^{(k)}_{t}:=1+\sum_{n=1}^{t}\sqrt{\varphi^{(k)}_{n}}$. Then for all $t\geq 1$,
$\mathbb{P}\left(L^{(k^{\star})}_{t}>\min_{k\neq k^{\star}}L^{(k)}_{t}\right)$ $\leq 2\sum_{k\neq k^{\star}}\exp\left(-\frac{t\Delta_{k}^{2}}{8\max\{(\theta_{t}^{(k)})^{2},(\theta_{t}^{(k^{\star})})^{2}\}(8\max\{v_{k^{\star}},v_{k}\}+\Delta_{k})}\right).$ (25)
If Assumption 2(ii) holds and for each $k\in[K]$, $\{\ell_{t}^{(k)}\}_{t}$ is geometrically $\varphi$-mixing with rate $\varphi^{(k)}_{n}\leq c_{k}\exp{(-b_{k}n^{\gamma_{k}})}$ with $b_{k}>0$, $c_{k}\geq 0$, and $\gamma_{k}>0$. Define $\gamma_{\min}:=\min_{k\in[K]}\gamma_{k}$, $b_{\min}:=\min_{k\in[K]}b_{k}$, $c_{\max}:=\max_{k\in[K]}c_{k}$, and
$t_{0}:=\max\big\{\min\left\{m\geq 3\>|\>m^{2}\geq 3232\,c_{\max}\textnormal{ and }m(\log m)^{-2/\gamma_{\min}}\geq 4\right\},e^{3/b_{\min}}\big\}.$ (26)
Then for all $t\geq t_{0}$
$\mathbb{P}$ $\left(L^{(k^{\star})}_{t}>\min_{k\neq k^{\star}}L^{(k)}_{t}\right)$
$\leq 4\sum_{k\neq k^{\star}}\exp\left(-\frac{t\Delta_{k}^{2}}{32(\max\left\{(\log t)^{2/{\gamma_{k^{\star}}}}v_{k^{\star}},(\log t)^{2/{\gamma_{k}}}v_{k}\right\}+\frac{1}{6}(\log t)^{2/\min\{\gamma_{k^{\star}},\gamma_{k}\}}\Delta_{k})}\right).$ (27)
We start by defining an event of interest $\mathcal{A}:=\{L_{t}^{(k^{\star})}>L_{t}^{(k)}\}$, $k\in[K]$. Notice that $L_{t}^{(k^{\star})}-L_{t}^{(k)}\geq 0$ and hence for any $t>0$ it holds that
$L_{t}^{(k^{\star})}-L_{t}^{(k)}+t\Delta_{k}-t\Delta_{k}=(L_{t}^{(k^{\star})}-t\mu_{k^{\star}})-(L_{t}^{(k)}-t\mu_{k})\geq t\Delta_{k}.$ (28)
Next, notice that, by Remark 2.1, one can define $S_{t}^{(k)}:=L_{t}^{(k)}-t\mu_{k}=\sum_{\tau=1}^{t}\left(\ell_{\tau}^{(k)}-\mu_{k}\right)$ for all $t\geq 1$ and $k\in[K]$, which allows us to write (28) as
$\mathcal{A}=\{S_{t}^{(k^{\star})}-S_{t}^{(k)}\geq t\Delta_{k}\}\subseteq\{S_{t}^{(k^{\star})}\geq\tfrac{1}{2}t\Delta_{k}\}\cup\{-S_{t}^{(k)}\geq\tfrac{1}{2}t\Delta_{k}\}$
and hence
$\mathbb{P}(\mathcal{A})\leq\mathbb{P}(S_{t}^{(k^{\star})}\geq\tfrac{1}{2}t\Delta_{k})+\mathbb{P}(S_{t}^{(k)}\leq-\tfrac{1}{2}t\Delta_{k}).$ (29)
Case (i). We first consider the case where Assumption 2(i) holds. Together with Assumption 1, it implies that $\{S_{t}^{(k)}\}_{t}$ are sums of bounded zero-mean random variables with variances $\{v^{(k)}\}_{k}$. We use the one-sided Bernstein inequality (see Lemma A.5) and obtain
$\mathbb{P}(S_{t}^{(k^{\star})}\geq\tfrac{1}{2}t\Delta_{k})\leq\exp\left(-\frac{t\Delta_{k}^{2}}{8v_{k^{\star}}+\tfrac{4}{3}\Delta_{k}}\right)\kern 5.0pt\text{and}\kern 5.0pt\mathbb{P}(S_{t}^{(k)}\leq-\tfrac{1}{2}t\Delta_{k})\leq\exp\left(-\frac{t\Delta_{k}^{2}}{8v_{k}+\tfrac{4}{3}\Delta_{k}}\right).$
Substituting these bounds in (29) and using the elementary inequality in Lemma A.1, one obtains that
$\mathbb{P}(\mathcal{A})\leq 2\exp\left(-\frac{t\Delta_{k}^{2}}{8\max\{v_{k^{\star}},v_{k}\}+\tfrac{4}{3}\Delta_{k}}\right),$
which, by the union bound, yields the inequality in (24).
Case (ii). We proceed analogously to the proof of case (i), using a one-sided Bernstein inequality in Lemma A.6 for sums of $\varphi$-mixing random variables in (29). Using that, by Assumption 1, $M=1$ and choosing $\varepsilon=\frac{1}{2}\Delta_{k}t$ in (20) and (21) yields the following bounds
$\mathbb{P}(S_{t}^{(k^{\star})}\geq\tfrac{1}{2}t\Delta_{k})$ $\leq\exp\left(-\frac{t\Delta_{k}^{2}}{8(\theta_{t}^{(k^{\star})})^{2}(8v_{k^{\star}}+\Delta_{k})}\right),$
$\mathbb{P}(S_{t}^{(k)}\leq-\tfrac{1}{2}t\Delta_{k})$ $\leq\exp\left(-\frac{t\Delta_{k}^{2}}{8(\theta_{t}^{(k)})^{2}(8v_{k}+\Delta_{k})}\right).$
Using these inequalities in (29) together with the elementary result in Lemma A.1 results in
$\mathbb{P}(\mathcal{A})\leq 2\exp\left(-\frac{t\Delta_{k}^{2}}{8\max\{(\theta_{t}^{(k)})^{2},(\theta_{t}^{(k^{\star})})^{2}\}(8\max\{v_{k^{\star}},v_{k}\}+\Delta_{k})}\right),$
which, by the union bound, yields the inequality in (25) for all $t\geq 1$.
Case (iii). The proof is analogous to the proof of case (i), but now using a one-sided Bernstein inequality for sums of geometrically $\varphi$-mixing random variables in (29), see Theorem A.1 in Hang, (2015); Hang and Steinwart, (2017). Taking $\varepsilon=\frac{1}{2}\Delta_{k}t$ yields, for all $t\geq t_{0}$ with $t_{0}$ as in (26), obtained as the largest among all the experts, the following one-sided bounds:
$\mathbb{P}\left(S_{t}^{(k^{\star})}\geq\tfrac{t\Delta_{k}}{2}\right)$ $\leq 2\exp\left(-\frac{t^{2}\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k^{\star}}}(v_{k^{\star}}+\tfrac{1}{6}t\Delta_{k})}\right)\leq 2\exp\left(-\frac{t\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k^{\star}}}(v_{k^{\star}}+\tfrac{1}{6}\Delta_{k})}\right),$
$\mathbb{P}\left(S_{t}^{(k)}\leq-\tfrac{t\Delta_{k}}{2}\right)$ $\leq 2\exp\left(-\frac{t^{2}\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k}}(v_{k}+\tfrac{1}{6}t\Delta_{k})}\right)\leq 2\exp\left(-\frac{t\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k}}(v_{k}+\tfrac{1}{6}\Delta_{k})}\right),$
where we used that $M=1$ by Assumption 1, and that ${t^{2}}/(a+bt)>t/(a+b)$ for any $a,b,t>0$. Substituting these bounds in (29) and using the fact that $\exp(-\tfrac{x}{a})+\exp(-\tfrac{x}{b})\leq 2\exp(-\tfrac{x}{\max\{a,b\}})$ for $x\geq 0$ and $a,b>0$, one obtains that
$\mathbb{P}(\mathcal{A})\leq 4\exp\left(-\frac{t\Delta_{k}^{2}}{32(\max\left\{(\log t)^{2/{\gamma_{k^{\star}}}}v_{k^{\star}},(\log t)^{2/{\gamma_{k}}}v_{k}\right\}+\frac{1}{6}(\log t)^{2/\min\{\gamma_{k^{\star}},\gamma_{k}\}}\Delta_{k})}\right),$
which, by the union bound, yields (27). ∎
Corollary B.2.
Let the conditions of Lemma B.2 hold and let $v_{\max}:=\max_{k\in[K]}v_{k}$. Then the following bounds hold:
Under the conditions of case (i) of Lemma B.2
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq 2(K-1)\exp\left(-\frac{t\Delta^{2}}{8v_{\max}+\tfrac{4}{3}\Delta}\right).$ (30)
Under the conditions of case (ii) of Lemma B.2, assume, additionally, that Assumption 4(i) holds and define $\overline{\theta}_{\max}^{2}=\max_{k\in[K]}\sup_{t\geq 1}(\theta^{(k)}_{t})^{2}$. Then
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq{2}(K-1)\exp\left(-\frac{t\Delta^{2}}{8\overline{\theta}_{\max}^{2}}(8v_{\max}+\Delta)\right).$ (31)
Under the conditions of case (iii) of Lemma B.2,
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq 4(K-1)\exp\left(-\frac{t(\log t)^{-2/\gamma_{\min}}\Delta^{2}}{32(v_{\max}+\tfrac{1}{6}\Delta)}\right).$ (32)
Part (i) is an immediate consequence of case (i) of Lemma B.2. Using the elementary inequality in Lemma A.1, and noticing that the exponent of the exponential is a decreasing function in $\Delta_{k}$ yields the required result. Part (ii) directly follows from case (i) of Lemma B.2, using again that the exponent of the exponential is a decreasing function in $\Delta_{k}$ as well as in $(\theta^{(k)}_{t})^{2}$. Additionally, by Assumption 4(i), $\overline{\theta}_{\max}^{2}=\max_{k\in[K]}\sup_{t\geq 1}\left(\theta_{t}^{(k)}\right)^{2}\leq\left(1+C_{2,\varphi}\right)^{2}<\infty$. Part (iii) directly follows from case (iii) of Lemma B.2 by noticing that it holds that $\Delta_{k}\geq\Delta$ and
$\max\{(\log t)^{2/\gamma_{k^{\star}}}v_{k^{\star}},(\log t)^{2/\gamma_{k}}v_{k}\}\leq(\log t)^{2/\gamma_{\min}}\max_{k\in[K]}v_{k}\quad{\rm for}\kern 5.0pt{\rm all}\kern 5.0ptk\in[K].$
B.1.3 Combined Bound
The following proposition provides upper bounds for the probability that the expert with the lowest expected loss commits a cumulative error that is larger than that of the runner-up expert. We combine both Hoeffding-type and Bernstein-type results to showcase the different kinds of guarantees that can be obtained. Note that, although the presented bounds assume constant-in-time second moment of losses, they remain valid if one uses time-dependent variances or their uniform-in-time bound.
Proposition B.1.
Let Assumptions 1-2 hold and, without loss of generality, assume that the experts are ordered such that $\mu_{1}<\mu_{2}\leq\ldots\leq\mu_{K}$, that is $\Delta=\mu_{2}-\mu_{1}>0$. Additionally, let $v_{k}:=\textnormal{Var}(\ell^{(k)}_{t})$, $k\in[K]$. Then for all $t\geq t_{0}$,
$\mathbb{P}\left(L^{(1)}_{t}>\min_{2\leq k\leq K}L^{(k)}_{t}\right)\leq\min\left\{H(t),B(t)\right\}$
with the following cases:
If the dependence conditions of Lemma B.2(i) hold, then then
$H(t)$ $=\exp\left(-\frac{t\Delta^{2}}{2}\right)\left[2+\sum_{k=3}^{K}\exp\left(-2t(\mu_{1}-\mu_{k})(\mu_{2}-\mu_{k})\right)\right],$
$B(t)$ $=2\sum_{k=2}^{K}\exp\left(-\frac{t\Delta_{k}^{2}}{8\max\{v_{1},v_{k}\}+\tfrac{4}{3}\Delta_{k}}\right),$
where $t_{0}=1$.
If the dependence conditions of Lemma B.2(ii) hold, then
$H(t)$ $=\sqrt{e}\sum_{k=1}^{K}\exp\left(-\frac{t(\mu_{1}-\mu_{k}+\Delta/2)^{2}}{2\theta^{(k)}_{t}}\right),\kern 5.0pt\text{with}\kern 5.0pt\theta^{(k)}_{t}=1+4\sum_{n=1}^{t-1}\varphi^{(k)}_{n}$,
$B(t)$ $={2}\sum_{k=2}^{K}\exp\left(-\frac{t\Delta_{k}^{2}}{8\,\max\{(\widetilde{\theta}^{(1)}_{t})^{2},\widetilde{\theta}^{(k)}_{t})^{2}\}(8\,\max\{v_{1},v_{k}\}+\Delta_{k})}\right),\kern 5.0pt\text{with}\kern 5.0pt\widetilde{\theta}^{(k)}_{t}=1+\sum_{n=1}^{t}\sqrt{\varphi^{(k)}_{n}}$.
If the dependence conditions of Lemma B.2(iii) hold, then
$H(t)$ $=\sqrt{e}\sum_{k=1}^{K}\exp\left(-\frac{t(\mu_{1}-\mu_{k}+\Delta/2)^{2}}{2\theta^{(k)}_{t}}\right),\kern 5.0pt\text{with}\kern 5.0pt\theta^{(k)}_{t}:=1+4\sum_{n=1}^{t-1}\varphi^{(k)}_{n}$,
$B(t)$ $=4\sum_{k=2}^{K}\exp\left(-\frac{t\Delta_{k}^{2}}{32(\max\left\{(\log t)^{2/{\gamma_{1}}}v_{1},(\log t)^{2/{\gamma_{k}}}v_{k}\right\}+\frac{1}{6}(\log t)^{2/\min\{\gamma_{1},\gamma_{k}\}}\Delta_{k})}\right),$
where $t_{0}$ is given explicitly in (26), Lemma B.2, and $\gamma_{k}>0$ for all $k\in[K]$.
The assumptions allow us to apply both Lemma B.1 and B.2.
Case (i). Take $H(t)$ to be the Hoeffding-type bound from Lemma B.1(i) and $B(t)$ the Bernstein-type bound from Lemma B.2(i).
Case (ii). Under the $\varphi$-mixing assumptions, take $H(t)$ to be the Hoeffding-type bound from Lemma B.1(ii) and $B(t)$ the Bernstein-type bound from Lemma B.2(ii).
Case (iii). Under geometrically $\varphi$-mixing assumptions, similarly take $H(t)$ to be the bound from Lemma B.1(ii) and $B(t)$ from Lemma B.2(iii).
In either case, when $t\geq t_{0}$, we can control the probability of $L^{(1)}_{t}$ being greater than all other cumulative expert losses by the minimum of the two bounds, since both apply. ∎
B.2 Concentration Results for Hedge
We exploit the theory from Appendix A.2 to construct concentration-type bounds for the cumulative losses of experts under both i.i.d. and dependent settings.
B.2.1 Hoeffding-type
Lemma B.3 (Hoeffding-type).
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss among $K$ experts. Then for $k\in[K]$ and any $t$
If Assumption 2(i) holds, then
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\exp\left(-\frac{\Delta_{k}^{2}t}{8}\right).$ (33)
If Assumption 2(iii) holds, then
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\sqrt{e}\exp\left(-\frac{\Delta_{k}^{2}t}{8{\rho^{(k)}_{t}}}\right)$, (34)
where $\rho^{(k)}_{t}:=1+\sum_{s=1}^{t}\varphi_{s}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)$.
In order to derive the bound for $\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<{\Delta_{k}t}/{2}\right)$, one constructs, for every $k\neq k^{\star}$, random variables $Z_{t}^{(k)}:=-\ell_{t}^{(k)}+\ell_{t}^{(k^{\star})}+\Delta_{k}$, $Z_{t}^{(k)}\in[-1+\Delta_{k},1+\Delta_{k}]\subset[-1,2]$, $\mathbb{E}[Z_{t}^{(k)}]=0$, and writes
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $=\mathbb{P}\left(\sum_{s=1}^{t}Z_{s}^{(k)}>\frac{\Delta_{k}t}{2}\right).$ (35)
We now apply the conditions in case (i) and case (ii) to conclude the proof.
Case (i). Under Assumption 2(i), for each $k\in[K]$, $\{Z_{t}^{(k)}\}_{t}$ are i.i.d. and Hoeffding’s inequality yields
$\mathbb{P}\left(\sum_{s=1}^{t}Z_{s}^{(k)}>\frac{\Delta_{k}t}{2}\right)\leq\exp\left({-\frac{t}{2}\left(\frac{\Delta_{k}}{2}\right)^{2}}\right)\leq\exp\left(-\frac{\Delta_{k}^{2}t}{8}\right),$
which is the inequality in (33).
Case (ii). Under Assumption 2(iii), for each $k\in[K]$, $\{Z_{t}^{(k)}\}_{t}$ is $\varphi$-mixing (see, for example, Dedecker et al., (2007)) with the mixing coefficients bounded by the mixing coefficients $\{\varphi_{n}\}_{n\geq 1}$ of the $K$ experts’ losses sequence. Then by Lemma A.4 and definition of ${\rho}^{(k)}_{t}$, for all $t\geq 1$,
$\mathbb{P}\left(\sum_{s=1}^{t-1}Z^{(k)}_{s}>\frac{\Delta_{k}(t-1)}{2}\right)\leq\sqrt{e}\exp\left(-\frac{\Delta_{k}^{2}(t-1)}{8{\rho}^{(k)}_{t}}\right)$, (36)
which yields (34). ∎
Remark B.1.
Notice that given some bound
$\mathbb{P}\left(L^{(k)}_{t}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq a,$
for all $k\not=k^{*}$, one can use that
$\left\{\min_{k\neq k^{\star}}L_{t}^{(k)}-L_{t}^{\left(k^{\star}\right)}<\frac{\Delta t}{2}\right\}\subseteq\bigcup_{k\neq k^{\star}}\left\{L_{t}^{(k)}-L_{t}^{\left(k^{\star}\right)}<\frac{\Delta_{k}t}{2}\right\}$
to obtain
$\mathbb{P}\left(\min_{k\neq k^{\star}}L_{t}^{(k)}-L_{t}^{\left(k^{\star}\right)}<\frac{\Delta t}{2}\right)\leq(K-1)a.$
Corollary B.3.
Let the conditions of Lemma B.3 hold. Then the following bounds hold:
Under the conditions of case (i) of Lemma B.3
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)\leq\exp\left(-\frac{\Delta^{2}t}{8}\right).$ (37)
Under the conditions of case (ii) of Lemma B.3, and, additionally, under Assumption 3(ii)
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)\leq\sqrt{e}\exp\left(-\frac{\Delta^{2}t}{8\overline{\rho}_{\max}}\right),$ (38)
where $\overline{\rho}_{\max}=\max_{k\not=k^{\star}}\sup_{t\geq 1}\rho^{(k)}_{t}$.
To obtain case (i), we simply note that $\Delta_{k}\geq\Delta$ for all $k\in[K]$ on the right-hand side of (33). Case (ii) similarly follows by using that by Assumption 3(ii), $\overline{\rho}_{\max}\leq 1+\widetilde{C}_{1,\phi}<\infty$, and upper bounding (34). ∎
B.2.2 Bernstein-type
Lemma B.4 (Bernstein-type).
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss among $K$ experts. Define $\widetilde{v}_{k}:=\textnormal{Var}(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})$, $\in[K], kk$$\neq k^{\star}$. Then
If Assumption 2(i) holds, then
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\exp\left(-\frac{t\Delta_{k}^{2}}{8(\widetilde{v}_{k}+\frac{1}{3}\Delta_{k})}\right).$ (39)
If Assumption 2(iii) holds, then
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\exp\left(-\frac{t\Delta_{k}^{2}}{16({\rho}^{(k)}_{t})^{2}(4\widetilde{v}_{k}+\Delta_{k})}\right),$ (40)
where ${\rho}^{(k)}_{t}:=1+\sum_{n=1}^{t}\sqrt{\varphi_{n}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)}$.
If Assumption 2(iii) holds and for each $k\in[K]$, $\{\ell_{t}^{(k)}-\ell_{t}^{(k^{\star})}\}_{t}$ is geometrically $\varphi$-mixing with rate $\varphi^{(k)}_{n}\leq c_{k}\exp{(-b_{k}n^{\gamma_{k}})}$ with $b_{k}>0$, $c_{k}\geq 0$, and $\gamma_{k}>0$. Define $\gamma_{\min}:=\min_{k\in[K]}\gamma_{k}$, $b_{\min}:=\min_{k\in[K]}b_{k}$, $c_{\max}:=\max_{k\in[K]}c_{k}$, and
$t_{0}:=\max\big\{\min\left\{m\geq 3\>|\>m^{2}\geq 6464\,c_{\max}\textnormal{ and }m(\log m)^{-2/\gamma_{\min}}\geq 4\right\},e^{3/b_{\min}}\big\}.$ (41)
Then for all $t\geq t_{0}$
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq 2\exp\left(-\frac{t^{2}\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k}}\left(\widetilde{v}_{k}+\frac{t}{3}\Delta_{k}\right)}\right)$ (42)
Analogously to the proof of Lemma B.3, we separately analyze (35) for cases (i) and (ii).
Case (i). Under Assumption 2(i), $\{Z_{t}^{(k)}\}_{t}$ are i.i.d. for each $k\in[K]$. We apply the one-sided Bernstein inequality (see Lemma A.5) and obtain
$\mathbb{P}\left(\sum_{s=1}^{t}Z^{(k)}_{s}>\frac{\Delta_{k}t}{2}\right)\leq\exp\left(-\frac{\Delta_{k}^{2}t}{8(\widetilde{v}_{k}+\frac{1}{3}\Delta_{k})}\right),$ (43)
where we used that $M=2$, since $Z_{t}^{(k)}\in[-1+\Delta_{k},1+\Delta_{k}]\subset[-1,2]$. This expression is the inequality in (39), as required.
Case (ii). Under Assumption 2(iii), $\{Z_{t}^{(k)}\}_{t}$ are $\varphi$-mixing for each $k\in[K]$ with coefficients $\{\varphi_{n}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)\}_{n\geq 1}$. The absolute value of random variables $\{Z_{t}^{(k)}\}_{t}$ is bounded from above by $M=2$ by Assumption 1. We apply the one-sided Bernstein inequality (see Lemma A.6) and obtain
$\mathbb{P}\left(\sum_{s=1}^{t}Z^{(k)}_{s}>\frac{\Delta_{k}t}{2}\right)\leq\exp\left(-\frac{\Delta_{k}^{2}t}{16({\rho}^{(k)}_{t})^{2}(4\widetilde{v}_{k}+\Delta_{k})}\right),$ (44)
where we define ${\rho}^{(k)}_{t}$} for all $t\in[T]$ and $k\in[K]$ as in the statement, and (40) follows as required.
Case (iii). The proof is analogous to the proof of case (i) but now using a one-sided Bernstein inequality for sums of geometrically $\varphi$-mixing random variables $\{Z_{t}^{(k)}\}_{t}$, see Theorem A.1 (Hang,, 2015; Hang and Steinwart,, 2017). Taking $\varepsilon=\frac{1}{2}\Delta_{k}t$ yields, for all $t\geq t_{0}$ with $t_{0}$ as in (41), obtained as the largest among all the experts and taking into account that $M=2$, the following bound:
$\mathbb{P}\left(\sum_{s=1}^{t}Z^{(k)}_{s}>\frac{\Delta_{k}t}{2}\right)$ $\leq 2\exp\left(-\frac{\frac{1}{4}t^{2}\Delta_{k}^{2}}{8(\log t)^{2/\gamma_{k}}\left(\widetilde{v}_{k}+\frac{\Delta_{k}t}{3}\right)}\right)=2\exp\left(-\frac{t^{2}\Delta_{k}^{2}}{32(\log t)^{2/\gamma_{k}}\left(\widetilde{v}_{k}+\frac{\Delta_{k}t}{3}\right)}\right),$
which is (42). ∎
Corollary B.4.
Let the conditions of Lemma B.4 hold. Then the following bounds hold:
Under the conditions of case (i) of Lemma B.4
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\exp\left(-\frac{t\Delta^{2}}{8(\max_{k\neq k^{\star}}\widetilde{v}_{k}+\frac{1}{3}\Delta)}\right).$ (45)
Under the conditions of case (ii) of Lemma B.4, let, additionally, Assumption 4(ii) holds. Then
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq\exp\left(-\frac{t\Delta^{2}}{16\,\overline{\rho}_{\max}^{2}(4\max_{k\neq k^{\star}}\widetilde{v}_{k}+\Delta)}\right).$ (46)
where $\overline{\rho}_{\max}=\max_{k\in[K]}\sup_{t\geq 1}\rho^{(k)}_{t}$.
Under the conditions of case (iii) of Lemma B.4,
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)$ $\leq 2\exp\left(-\frac{t(\log t)^{-2/\gamma_{\min}}\Delta^{2}}{32(\max_{k\neq k^{\star}}\widetilde{v}_{k}+\frac{1}{3}\Delta)}\right)$ (47)
for all $t\geq t_{0}$.
Case (i) is obtained by using that $\Delta_{k}\geq\Delta$ for all $k\in[K]$, exactly as in the proof of Corollary B.2. For case (ii), one notes that on the right-hand side of (40) $\Delta_{k}\geq\Delta$ for all $k\in[K]$ and $\overline{\rho}_{\max}^{2}\geq(\rho_{t}^{(k)})^{2}$ for all $t\in[T]$ and $k\in[K]$. Notice now that, by Assumption 4(ii), $\overline{\rho}_{\max}^{2}\leq(1+\widetilde{C}_{2,\varphi})^{2}<\infty$. Lastly, to obtain the bound of case (iii), we additionally observe that
$(\log t)^{2/\gamma_{k}}\left(\widetilde{v}_{k}+\frac{t}{3}\Delta_{k}\right)$ $\leq(\log t)^{2/\gamma_{k}}t\left(\widetilde{v}_{k}+\frac{1}{3}\Delta_{k}\right)\leq(\log t)^{2/\gamma_{\rm min}}t\left(\max_{k\neq k^{\star}}\widetilde{v}_{k}+\frac{1}{3}\Delta\right),$
for all $k\in[K]$. ∎
B.2.3 Combined Bound
The following results contain the necessary concentration results together with the bound on the expected leader regret as a function of sub-optimality gap $\Delta$. We present the results for the i.i.d. setting as well as under the $\varphi$-mixing assumption on losses.
Proposition B.2.
Let Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge. Let $k^{\star}\in[K]$ be the expert with the smallest expected loss among $K$ experts. Define $\widetilde{v}_{k}:=\textnormal{Var}(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})$, $\in[K], kk$$\neq k^{\star}$. Then for all $t\geq t_{0}$,
$\mathbb{P}\left(L_{t}^{(k)}-L_{t}^{(k^{\star})}<\frac{\Delta_{k}t}{2}\right)\leq\min\left\{H(t),B(t)\right\}$
with the following cases:
If the dependence conditions of Lemma B.4(i) hold, then
$H(t)=\exp\left(-\frac{\Delta_{k}^{2}t}{8}\right),\quad B(t)=\exp\left(-\frac{\Delta_{k}^{2}t}{8(\widetilde{v}_{k}+\frac{1}{3}\Delta_{k})}\right),\kern 5.0pt{\it with}\kern 5.0ptt_{0}=1.$
If the dependence conditions of Lemma B.4(ii) hold, then
$H(t)=\sqrt{e}\exp\left(-\frac{\Delta_{k}^{2}t}{8{\rho^{(k)}_{t}}}\right),\quad B(t)=\exp\left(-\frac{\Delta_{k}^{2}t}{16(\widetilde{\rho}^{(k)}_{t})^{2}(4\widetilde{v}_{k}+\Delta_{k})}\right),$
where $\rho^{(k)}_{t}:=1+\sum_{s=1}^{t}\varphi_{s}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)$, and $\widetilde{\rho}^{(k)}_{t}:=1+\sum_{n=1}^{t}\sqrt{\varphi_{n}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)}$, and $t_{0}=1$.
If the dependence conditions of Lemma B.4(iii) hold, then
$H(t)=\sqrt{e}\exp\left(-\frac{\Delta_{k}^{2}t}{8{\rho^{(k)}_{t}}}\right),\quad B(t)=2\exp\left(-\frac{\Delta_{k}^{2}t^{2}}{32(\log t)^{2/\gamma_{k}}\left(\widetilde{v}_{k}+\frac{t}{3}\Delta_{k}\right)}\right),$
where $\rho^{(k)}_{t}:=1+\sum_{s=1}^{t}\varphi_{s}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)$, $t_{0}$ is given explicitly in (41) in Lemma B.4, and $\gamma_{k}>0$ for all $k\in[K]$.
The proof of Proposition B.1 is a direct combination of the concentration results of Appendix B.1. With this proposition at hand, we can extend the time-splitting approach used by Mourtada and Gaïffas, (2019) to studying the regret of FTL under dependence.
The proof is identical to the proof of Proposition B.2, where Lemma B.3 and Lemma B.4 are used instead of Lemma B.1 and Lemma B.2, respectively. ∎
B.3 FTL Regret Bounds
Proposition B.3 (Hoeffding-type regret bounds).
Under the same setting of Lemma B.1, suppose that almost surely there are no ties in the FTL weights for all $t\geq 1$.
If Assumption 2(i) holds, then
$\mathbb{E}[\overline{R}_{\textnormal{{ftl}},T}]\leq 2+\frac{2\log K+4}{\Delta^{2}}.$
Let Assumption 2(ii) and Assumption 3(i) hold, and define $\overline{\theta}_{\max}:=\max_{k\in[K]}\sup_{t\geq 1}\theta^{(k)}_{t}$. Then
$\mathbb{E}[\overline{R}_{\textnormal{{ftl}},T}]\leq 3+\frac{8\overline{\theta}_{\max}(\log K+4)}{\Delta^{2}}.$
The bounds are uniform in $T$.
Our proof proceeds by exploiting the time splitting arguments also used by Mourtada and Gaïffas, (2019) for studying decreasing Hedge. Informally, the key idea is that one may split $[T]$ into “warm-up” and “effective” periods. By choosing the splitting time, $t_{0}$, carefully, we can obtain bounds that depend only logarithmically on the number of experts.
Recall that for $t\in[T]$ we define $c_{t}=\mathds{1}\{\textnormal{{ftl}}_{t-1}\not=\textnormal{{ftl}}_{t}\}$ (the leader set changes at $t$). We also define $C_{t}$ as the total number of times the leader set changes up to time $t\in[T]$, that is $C_{t}=\sum_{\tau=1}^{t}c_{\tau}$, using the fact that almost surely there are no ties in determining the FTL expert during any round. Introduce now the “warm-up time” $t_{0}\geq 2$, so that
$C_{T}=\sum_{t=1}^{T}c_{t}\leq t_{0}+\sum_{t=t_{0}+1}^{T}c_{t},$ (48)
where for periods $1\leq t\leq t_{0}$ we use $c_{t}=\mathds{1}\{\textnormal{{ftl}}_{t-1}\neq\textnormal{{ftl}}_{t}\}\leq 1$. We now write
$c_{t}$ $=\mathds{1}\{\textnormal{{ftl}}_{t-1}=\{1\}\wedge\textnormal{{ftl}}_{t}\neq\{1\}\}+\mathds{1}\{\textnormal{{ftl}}_{t-1}\neq\{1\}\wedge\textnormal{{ftl}}_{t}=\{1\}\}$
$=\mathds{1}\{\textnormal{{ftl}}_{t-1}=\{1\}\}\cdot\mathds{1}\{\textnormal{{ftl}}_{t}\neq\{1\}\}+\mathds{1}\{\textnormal{{ftl}}_{t-1}\neq\{1\}\}\cdot\mathds{1}\{\textnormal{{ftl}}_{t-1}=\{1\}\}$
$\leq\mathds{1}\{\textnormal{{ftl}}_{t-1}\neq\{1\}\}+\mathds{1}\{\textnormal{{ftl}}_{t}\neq\{1\}\}.$ (49)
Case (i). Using (49) we write that
$\mathbb{E}[c_{t}]$ $\leq\mathbb{P}(\textnormal{{ftl}}_{t-1}\neq\{1\})+\mathbb{P}(\textnormal{{ftl}}_{t}\neq\{1\})\leq K\left(e^{-(t-1)\Delta^{2}/2}+e^{-t\Delta^{2}/2}\right),$
where in the last inequality we used the bound (22) in Corollary B.1.
Next, we note that for $t\geq t_{0}+1$
$K\left(e^{-(t-1)\Delta^{2}/2}+e^{-t\Delta^{2}/2}\right)$ $=\left(Ke^{-t_{0}\Delta^{2}/2}\right)e^{-(t-t_{0}-1)\Delta^{2}/2}+\left(Ke^{-t_{0}\Delta^{2}/2}\right)e^{-(t-t_{0})\Delta^{2}/2}$
$\leq e^{-(t-t_{0}-1)\Delta^{2}/2}+e^{-(t-t_{0})\Delta^{2}/2}.$ (50)
Finally, we choose $t_{0}=\lceil\frac{2\log K}{\Delta^{2}}\rceil$ using (22) in Corollary B.1, and, taking expectation on both sides of (48), write
$\mathbb{E}[C_{T}]\leq 1+\frac{2\log K}{\Delta^{2}}+\left(1+2\sum_{s=1}^{\infty}e^{-s\Delta^{2}/2}\right)\leq 2+\frac{2\log K}{\Delta^{2}}+\frac{4}{\Delta^{2}},$
where in the last inequality Lemma A.2 is applied. Finally, by Lemma 3.1 and noticing that $S_{T}\leq 1$, we obtain the following bound that does not depend on the terminal time $T$:
$\mathbb{E}[\overline{R}_{\textnormal{{ftl}},T}]\leq\mathbb{E}[C_{T}]\leq 2+\frac{2\log K+4}{\Delta^{2}},$
as required.
Case (ii). We start by noticing that by Assumption 3(i), $\overline{\theta}_{\max}=\max_{k\in[K]}\sup_{t\geq 1}\{1+4\sum_{n=1}^{t-1}\varphi_{n}^{(k)}\}\leq 1+4C_{1,\varphi}<\infty$. Next, similarly to case (i), using (49) we write that
$\mathbb{E}[c_{t}]$ $\leq\mathbb{P}(\textnormal{{ftl}}_{t-1}\neq\{1\})+\mathbb{P}(\textnormal{{ftl}}_{t}\neq\{1\})\leq K\sqrt{e}\left[\exp\left(-\frac{(t-1)\Delta^{2}}{8\overline{\theta}_{\max}}\right)+\exp\left(-\frac{t\Delta^{2}}{8\overline{\theta}_{\max}}\right)\right],$
where in the last inequality we used the bound (23) in Corollary B.1. The same bound allows to set $t_{0}=\lceil\frac{8\overline{\theta}_{\max}\log K}{\Delta^{2}}\rceil$. Hence, (48) together with Lemma A.2 yields
$\mathbb{E}[C_{T}]\leq 1+\frac{8\overline{\theta}_{\max}\log K}{\Delta^{2}}+\sqrt{e}\left(1+2\sum_{s=0}^{\infty}e^{-s\Delta^{2}/(8\overline{\theta}_{\max})}\right)\leq 3+\frac{8\overline{\theta}_{\max}\log K}{\Delta^{2}}+\frac{32\overline{\theta}_{\max}}{\Delta^{2}}.$
Finally, utilizing this bound in Lemma 3.1 with $S_{T}\leq 1$ concludes the proof. ∎
Remark B.2.
Proposition B.3 provides bounds dependent on the squared sub-optimality gap $\Delta^{2}$, which are less sharp than results for decreasing Hedge with i.i.d. data, where only a factor $1/\Delta$ is required (see Mourtada and Gaïffas,, 2019, Theorem 2). Still, Proposition B.3(i) can be easily strengthened. Focusing on the expected pseudo-regret of Follow-the-Leader,
$\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},T}]:=\mathbb{E}\big[\,\overline{L}_{T}-L^{(k^{\star})}_{T}\big$],
we find that
$\mathbb{E}[r^{(k^{\star})}_{t}]=\mathbb{E}\left[\sum_{j=1}^{K}\omega^{(k)}_{\textnormal{{ftl}},t}\left(\ell^{(j)}_{t}-\ell^{(k^{\star})}_{t}\right)\right]=\sum_{j=1}^{K}\mathbb{E}[\omega^{(k)}_{\textnormal{{ftl}},t}]\Delta_{j},$ (51)
where the last equality follows from sequential independence of losses. For simplicity, let us assume again a no-tie condition, meaning $\lvert\textnormal{{ftl}}_{t}\rvert=1$ for all $t\geq 1$. Using the same setting of Lemma B.1, we have that $\Delta_{1}=\mathbb{E}[\ell^{(1)}_{t}-\ell^{(k^{\star})}_{t}]=0$ and, for $2\leq k\leq K$, $\mathbb{E}[\omega^{(k)}_{\textnormal{{ftl}},t}]\leq\mathbb{P}(\textnormal{{ftl}}_{t}\not=\{1\})$, hence $\mathbb{E}[r^{(k^{\star})}_{t}]\leq\sum_{j=2}^{K}\Delta_{j}\exp(-t\Delta_{j}^{2}/2)$. By following the same arguments used in the proof of Theorem 2 of Mourtada and Gaïffas, (2019) and Proposition B.3, a bound of the form
$\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},T}]\leq 1+\frac{3}{\Delta}+\frac{2\log(K)}{\Delta^{2}}$
can be obtained. This can then be straightforwardly extended to a bound for expected regret $\mathbb{E}[\overline{R}_{\textnormal{{ftl}},T}]$, as in Remark 14 of Mourtada and Gaïffas, (2019). When working with dependent losses, however, the chain of equalities in (51) cannot be used, leading to the coarser result of Proposition B.3(ii).
Splitting the expected pseudo regret again at $t_{0}=\lceil\frac{2\log(K)}{\Delta^{2}}\rceil$, we find
$\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},T}]\leq\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},t_{0}}]+\sum_{t=t_{0}+1}^{T}\mathbb{E}[r^{(k^{\star})}_{t}$].
As before, a trivial bound on the worst-case regret until $t_{0}$ yields $\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},t_{0}}]\leq 1+2\log(K)/\Delta^{2}$. Since the mapping $x\mapsto xe^{-x^{2}/2}$ is decreasing over $1,\infty)$ and $\Delta_{j}\geq\Delta$, whenever $t\geq 1+1/\Delta^{2}$ (which holds for $t\geq t_{0}+1$) we get
$\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},T}]\leq\Delta e^{-(t-t_{0}-1)\Delta^{2}/2},$
where we have followed the derivation steps used for (B.3) and the fact that $K\exp(-t_{0}\Delta^{2}/2)\leq 1$. Applying Lemma A.2 and collecting terms,
$\mathbb{E}[{R}^{(k^{\star})}_{\textnormal{{ftl}},T}]\leq 1+\frac{2\log(K)}{\Delta^{2}}+\Delta\left(1+\frac{2}{\Delta^{2}}\right)\leq 1+\frac{3}{\Delta}+\frac{2\log(K)}{\Delta^{2}}.$
Proposition B.4 (Bernstein-type regret bounds).
Let the conditions of Lemma B.2 hold. Assume that there are almost surely no ties in the FTL weights for all $t\geq 1$, and define $v_{\max}:=\max_{k\in[K]}v_{k}$.
Under the conditions of case (i) of Lemma B.2 it holds
$\mathbb{E}\big[\overline{R}_{\textnormal{{ftl}},T}\big]\leq\ 2+\left(\log(2K)+2\right)\frac{8v_{\max}+\tfrac{4}{3}\Delta}{\Delta^{2}}.$
Under the conditions of case (ii) of Lemma B.2, and, additionally, under Assumption 4(i), define $\overline{\theta}_{\max}^{2}=\max_{k\in[K]}\sup_{t\geq 1}(\theta^{(k)}_{t})^{2}$. Then
$\mathbb{E}\big[\overline{R}_{\textnormal{{ftl}},T}\big]\leq\ 2+\left(\log(K)+2\right)\frac{8\overline{\theta}_{\max}^{2}(8v_{\max}+\Delta)}{\Delta^{2}}.$
Under the conditions of case (iii) of Lemma B.2, it holds
$\mathbb{E}\big[\overline{R}_{\textnormal{{ftl}},T}\big]$ $\leq\widetilde{t}_{0}+1+2\sum_{t=\widetilde{t}_{0}}^{T}\exp\left(-\frac{t\widetilde{c}_{\Delta}}{2(\log t)^{\alpha}}\right),$
where $\alpha=2/\gamma_{\rm min}$,
$\widetilde{c}_{\Delta}$ $=\frac{\Delta^{2}}{32(v_{\max}+\Delta/6)},$
and $\widetilde{t}_{0}$ is defined as
$\widetilde{t}_{0}=\left\lceil\inf\left\{t\geq\max\left\{{e^{\alpha}},{t_{0}}\right\}\,\bigg|\,\frac{t}{(\log t)^{\alpha}}\geq\frac{2\log(4K)}{\widetilde{c}_{\Delta}}\right\}\right\rceil,$
where $t_{0}$ is given in (26).
The bounds are uniform in $T$.
We proceed analogously to the proof strategy used for Proposition B.3, but applying instead the Bernstein-type results from Lemma B.2.
Case (i). By Corollary B.2(i)
$\mathbb{P}\left(\textnormal{{ftl}}_{t}\neq\{k^{\ast}\}\right)\leq 2(K-1)e^{-tc_{\Delta}}\leq 2Ke^{-tc_{\Delta}}\qquad\text{for all }t\geq 1,$ (52)
with $c_{\Delta}={\Delta^{2}}/{(8v_{\max}+\tfrac{4}{3}\Delta)}$. Recall that $C_{T}=\sum_{t=1}^{T}\mathds{1}\{\textnormal{{ftl}}_{t-1}\neq\textnormal{{ftl}}_{t}\}$, and let $t_{0}$ be a warm-up time so that, using (52), here too one obtains
$\mathbb{E}[C_{T}]$ $\leq t_{0}+\sum_{t=t_{0}+1}^{T}\big(\mathbb{P}(\textnormal{{ftl}}_{t-1}\neq\{k^{\ast}\})+\mathbb{P}(\textnormal{{ftl}}_{t}\neq\{k^{\ast}\})\big)$
$\leq t_{0}+2K\sum_{t=t_{0}+1}^{T}\left(e^{-(t-1)c_{\Delta}}+e^{-tc_{\Delta}}\right)$
$\leq t_{0}+2Ke^{-t_{0}c_{\Delta}}\sum_{t=t_{0}+1}^{T}\left(e^{-(t-t_{0}-1)c_{\Delta}}+e^{-(t-t_{0})c_{\Delta}}\right).$
Now, let the warm-up time be $t_{0}=\left\lceil\frac{\log(2K)}{c_{\Delta}}\right\rceil$, so that
$\mathbb{E}[C_{T}]\leq t_{0}+1+2\sum_{s=1}^{\infty}e^{-sc_{\Delta}}=t_{0}+1+\frac{2e^{-c_{\Delta}}}{1-e^{-c_{\Delta}}}\leq t_{0}+1+\frac{2}{c_{\Delta}}.$
Finally, $t_{0}\leq\frac{\log(2K)}{c_{\Delta}}+1$ gives
$\mathbb{E}[C_{T}]\leq\frac{\log(2K)}{c_{\Delta}}+2+\frac{2}{c_{\Delta}}=2+\left(\log(2K)+2\right)\frac{1}{c_{\Delta}}=2+\left(\log(2K)+2\right)\frac{8v_{\max}+\tfrac{4}{3}\Delta}{\Delta^{2}}.$
Since $\overline{R}_{\textnormal{{ftl}},T}\leq C_{T}$ under bounded losses ($S_{T}=1$), the stated bound follows.
Case (ii). We start by noticing that by Assumption 4(i), $\overline{\theta}_{\max}^{2}=\max_{k\in[K]}\sup_{t\geq 1}\{(1+\sum_{n=1}^{t-1}\sqrt{\varphi_{n}^{(k)}})^{2}\}\leq(1+C_{2,\varphi})^{2}<\infty$. Next, by Corollary B.2(ii), for $t\geq 1$ it holds
$\mathbb{P}\left(\textnormal{{ftl}}_{t}\neq\{k^{\ast}\}\right)\leq(K-1)e^{-{t\,\widetilde{c}_{\Delta}}}\leq Ke^{-{t\,\widetilde{c}_{\Delta}}},$
where
$\widetilde{c}_{\Delta}=\frac{\Delta^{2}}{8\overline{\theta}_{\max}^{2}(8v_{\max}+\Delta)}.$
Analogously to case (i), let the warm-up time be $t_{0}=\left\lceil\frac{\log(K)}{\widetilde{c}_{\Delta}}\right\rceil$, then
$\mathbb{E}[C_{T}]$ $\leq t_{0}+\sum_{t=t_{0}+1}^{T}\left(\mathbb{P}(\textnormal{{ftl}}_{t-1}\neq\{k^{\ast}\})+\mathbb{P}(\textnormal{{ftl}}_{t}\neq\{k^{\ast}\})\right)$
$\leq t_{0}+K\sum_{t=t_{0}+1}^{T}\left(e^{-(t-1)\widetilde{c}_{\Delta}}+e^{-t\widetilde{c}_{\Delta}}\right)$
$\leq t_{0}+Ke^{-t_{0}\widetilde{c}_{\Delta}}\sum_{t=t_{0}+1}^{T}\left(e^{-(t-t_{0}-1)\widetilde{c}_{\Delta}}+e^{-(t-t_{0})\widetilde{c}_{\Delta}}\right)$
$\leq\left(\frac{\log(K)}{\widetilde{c}_{\Delta}}+1\right)+\left(1+2\sum_{s=1}^{\infty}e^{-s\widetilde{c}_{\Delta}}\right)$
$\leq\frac{\log(K)}{\widetilde{c}_{\Delta}}+2+\frac{2}{\widetilde{c}_{\Delta}}=2+\left(\log(K)+2\right)\frac{8\overline{\theta}_{\max}^{2}(8v_{\max}+\Delta)}{\Delta^{2}},$
which concludes the proof.
Case (iii). By (27) in case (iii) of Lemma B.2, for $t\geq t_{0}$, with $t_{0}$ in (26) (and also from the simplified bound in (32)), it holds
$\mathbb{P}\left(\textnormal{{ftl}}_{t}\neq\{k^{\ast}\}\right)\leq 4K\exp\left(-\frac{t\,\widetilde{c}_{\Delta}}{(\log t)^{\alpha}}\right),$
where
$\widetilde{c}_{\Delta}=\frac{\Delta^{2}}{32(v_{\max}+\Delta/6)}$
and $\alpha=2/\gamma_{\min}$ for notational convenience. We use Lemma A.3 with $b=2\log(4K)/\widetilde{c}_{\Delta}$, which implies that there exists $\nu^{*}>0$ such that with $t^{*}:=\nu^{*}b$ one has $\frac{t^{*}}{\left(\log t^{*}\right)^{\alpha}}\geq b$. We now define the warm-up time as $\widetilde{t}_{0}:=\left\lceil\max\left\{t_{0},e^{\alpha},t^{*}\right\}\right\rceil$ with $t_{0}$ given in (26). Then, for all $t\geq\widetilde{t}_{0}$ it holds that
$\frac{t}{(\log t)^{\alpha}}\geq\frac{2\log(4K)}{\widetilde{c}_{\Delta}}$
and hence
$4K\exp\left(-\frac{t\,\widetilde{c}_{\Delta}}{(\log t)^{\alpha}}\right)$ $=\exp\left(\log(4K)-\frac{t\,\widetilde{c}_{\Delta}}{(\log t)^{\alpha}}\right)$
$\leq\exp\left(\frac{t\,\widetilde{c}_{\Delta}}{2(\log t)^{\alpha}}-\frac{t\,\widetilde{c}_{\Delta}}{(\log t)^{\alpha}}\right)\leq\exp\left(-\frac{t\,\widetilde{c}_{\Delta}}{2(\log t)^{\alpha}}\right).$
Proceeding as in cases (i) and (ii), we obtain
$\mathbb{E}[C_{T}]$ $\leq\widetilde{t}_{0}+1+4K\sum_{t=\widetilde{t}_{0}+1}^{T}\left\{\exp\left(-\frac{(t-1)\widetilde{c}_{\Delta}}{(\log(t-1))^{\alpha}}\right)+\exp\left(-\frac{t\widetilde{c}_{\Delta}}{(\log t)^{\alpha}}\right)\right\}$
$\leq\widetilde{t}_{0}+1+2\sum_{t=\widetilde{t}_{0}}^{T}\exp\left(-\frac{t\widetilde{c}_{\Delta}}{2(\log t)^{\alpha}}\right).$
Note that the series in the last line converges. Indeed, for any $C>0$ there exists $C^{\prime}>0$ such that $\exp(-sC/(\log s)^{\alpha})\lesssim\exp(-s^{C^{\prime}})$. Hence, it holds that $\sum_{t=\widetilde{t}_{0}}^{T}\exp\left(-\frac{t\widetilde{c}_{\Delta}}{2(\log t)^{\alpha}}\right)\leq\int_{0}^{\infty}\exp(-s^{C^{\prime}})\textnormal{d}s={\Gamma}(1/{C^{\prime}})/{C^{\prime}}<\infty$, with $\Gamma$} the Gamma function. Finally, as $\mathbb{E}\left[\overline{R}_{\textnormal{{ftl}},T}\right]\leq\mathbb{E}[C_{T}]$, one obtains a uniform in $T$ bound, as required. ∎
B.3.1 Combined Bounds for FTL
We can now obtain Theorem 3.1 by incorporating the bounds previously derived.
Part (i) follows immediately by combining Proposition B.3(i) and Proposition B.4(i). Similarly, part (ii) is obtained by combining Proposition B.3(ii) and Proposition B.4(ii), with the relevant constants and $t_{0}$ defined in the latter. ∎
B.4 Hedge Regret Bounds
Proposition B.5 (Hoeffding-type regret bounds).
Let $K\geq 3$ and Assumptions 1-2 hold. Let $\eta_{t}=2\sqrt{\log(K)/t}$ for $t\in[T]$ be the learning rate for decreasing Hedge.
If Assumption 2(i) holds,
$\mathbb{E}[\overline{R}_{\textnormal{{hdg}},T}]\leq\frac{4\Delta\log(K)+25}{\Delta^{2}}.$
If Assumption 2(ii) holds,
$\mathbb{E}[\overline{R}_{\textnormal{{hdg}},T}]\leq 2+\frac{(1+3\overline{\rho}_{\rm max})(\Delta\log(K)+16)}{\Delta^{2}},$
where $\overline{\rho}_{\rm max}:=\max_{k\not=k^{\star}}\sup_{t\geq 1}\rho^{(k)}_{t}$ with $\rho^{(k)}_{t}:=1+\sum_{s=1}^{t}\varphi_{s}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)$, $k\in[K]$.
As in the proof of Proposition B.3, time is split using some $t_{0}$ so that
$\overline{R}_{\textnormal{{hdg}},T}=\sum_{t=1}^{T}r_{t}=\overline{R}_{\textnormal{{hdg}},t_{0}}+\sum_{t=t_{0}+1}^{T}r_{t},$ (53)
where $\overline{R}_{\textnormal{{hdg}},t_{0}}$ is the average regret up to $t_{0}$. Using the worst-case bound (7) in Proposition A.1 in Chernov and Zhdanov, (2010) yields $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{t_{0}\log(K)}$. Then, for any $t\in\{t_{0}+1,\ldots,T\}$, we study the instantaneous expected Hedge regret
$\mathbb{E}[r_{t}]=\mathbb{E}\left[\sum_{k=1}^{K}\omega^{(k)}_{\textnormal{{hdg}},t}(\ell^{(k)}_{t}-\ell^{\dagger}_{t})\right]\leq\mathbb{E}\left[\sum_{k=1}^{K}\omega^{(k)}_{\textnormal{{hdg}},t}\left\lvert\ell^{(k)}_{t}-\ell^{\dagger}_{t}\right\rvert\right]\leq\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$],
where $\ell^{\dagger}_{t}:=\min_{k\in[K]}L^{(k)}_{t}$ is the minimal cumulative loss at time $t$ and is the loss of the empirical best expert at time $t$ (which, in general, is not the same as the best instantaneous $\ell^{*}_{t}$). We have used the fact that $\sum_{k=1}^{K}\omega^{(k)}_{\textnormal{{hdg}},t}=1$ for all $t\geq 1$, and that for any $t\in[T]$ it holds $\max_{k,k^{\prime}\in[K]}\lvert\ell^{(k)}_{t}-\ell^{(k^{\prime})}_{t}\rvert\leq 1$.
We now proceed to bound $\mathbb{E}[\omega^{(k)}_{\textnormal{{hdg}},t}$]. We mimic the steps in the proof of Theorem 2 in Mourtada and Gaïffas, (2019). More explicitly, we first notice that if $L_{t-1}^{(k)}-L_{t-1}^{(k^{\star})}\geq\frac{\Delta_{k}(t-1)}{2}$ then with $\eta_{t}=2\sqrt{\log(K)/t}$ one obtains
$\omega^{(k)}_{\textnormal{{hdg}},t}=\frac{\exp(-\eta_{t}(L^{(k)}_{t-1}-L^{(k^{\star})}_{t-1}))}{1+\sum_{j\neq k^{\star}}\exp(-\eta_{t}(L^{(j)}_{t-1}-L^{(k^{\star})}_{t-1}))}$ $\leq\exp\left(-2\sqrt{\frac{\log(K)}{t}}\frac{\Delta_{k}(t-1)}{2}\right)$
$\leq\exp\left(-\Delta_{k}\sqrt{\frac{(t-1)\log(K)}{2}}\right).$
Since $\omega^{(k)}_{\textnormal{{hdg}},t}\leq 1$, one can write, for $t\geq 2$,
$\omega^{(k)}_{\textnormal{{hdg}},t}=\mathds{1}\{L_{t-1}^{(k)}-L_{t-1}^{(k^{\star})}<{\Delta_{k}(t-1)}/{2}\}+\exp\left(-\Delta_{k}\sqrt{\frac{(t-1)\log(K)}{2}}\right),$
and, taking the expectation, one has
$\mathbb{E}[\omega^{(k)}_{\textnormal{{hdg}},t}]=\mathbb{P}\left(L_{t-1}^{(k)}-L_{t-1}^{(k^{\star})}<{\Delta_{k}(t-1)}/{2}\right)+\exp\left(-\Delta_{k}\sqrt{\frac{(t-1)\log(K)}{2}}\right).$ (54)
We now bound the first summand under Assumptions 2(i) and (ii) separately, which allows to complete the proof.
Case (i). Set $t_{0}=\lceil\frac{8\log(K)}{\Delta^{2}}\rceil$, which is chosen using (33) together with Remark B.1. Using case (i) of Corollary B.3 (see also the proof of Theorem 2 in Mourtada and Gaïffas, (2019)), we immediately note that, for $t\geq t_{0}+1$,
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$] $\leq(K-1)e^{-(t-1)\Delta^{2}/8}+(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/2}\,}$
$\leq((K-1)e^{-t_{0}\Delta^{2}/8})e^{-(t-t_{0}-1)\Delta^{2}/8}+((K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}})e^{-\Delta\sqrt{(t-1)\log(K)/8}}$
$\leq e^{-(t-t_{0}-1)\Delta^{2}/8}+e^{-\Delta\sqrt{(t-1)\log(K)/8}},$
where in the last inequality we used the fact that the choice of $t_{0}\geq{8\log(K)}/{\Delta^{2}}$ implies that $(K-1)e^{-t_{0}\Delta^{2}/8}\leq 1$. Then, for $t\geq t_{0}+1$ it holds that $(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\leq 1$. Hence,
$\sum_{t=t_{0}+1}^{\infty}\mathbb{E}[r_{t}]\leq\sum_{s=0}^{\infty}e^{-s\Delta^{2}/8}+\sum_{s=1}^{\infty}e^{-(\Delta/\sqrt{8})\sqrt{t}}\leq\dfrac{25}{\Delta^{2}},$ (55)
where we applied Lemma A.2 since $\log(K)>1$ (as we assumed $K\geq 3$).
We now consider again the bound that $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{t_{0}\log K}$, and, using our choice of $t_{0}$, we obtain $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq{\sqrt{8}\log K}/{\Delta}$. Therefore, in (53) we obtain
$\mathbb{E}\left[\overline{R}_{\textnormal{{hdg}},T}\right]\leq\sqrt{\log(K)}+\frac{\sqrt{8}\log(K)}{\Delta}+\frac{25}{\Delta^{2}}\leq\frac{4\Delta\log(K)+25}{\Delta^{2}},$
which is the required bound.
Case (ii). Set $t_{0}=\lceil\frac{8\overline{\rho}\log(K)}{\Delta^{2}}\rceil$, which is chosen using (34) together with Remark B.1. Case (ii) of Corollary B.3 implies that for $t\geq t_{0}+1$ it holds that
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$] $\leq\sqrt{e}\left((K-1)e^{-t_{0}\Delta^{2}/(8\overline{\rho})}\right)e^{-(t-t_{0}-1)\Delta^{2}/(8\overline{\rho})}$
$\quad\quad\quad+\left((K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\right)e^{-\Delta\sqrt{(t-1)\log(K)/8}}$
$\leq 2e^{-(t-t_{0}-1)\Delta^{2}/(8\overline{\rho})}+e^{-\Delta\sqrt{(t-1)\log(K)/8}},$
where the last inequality follows by noting that $\sqrt{e}\leq 2$ and $t_{0}\geq 8\overline{\rho}\log(K)/\Delta^{2}$ implies $(K-1)e^{-t_{0}\Delta^{2}/(8\overline{\rho})}\leq 1$, while $t\geq t_{0}+1$ also yields that $(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\leq 1$ since $\overline{\rho}\geq 1$ by definition. Therefore, again thanks to the assumption that $K\geq 3$ and by Lemma A.2,
$\sum_{t=t_{0}+1}^{T}\mathbb{E}[r_{t}]$ $\leq 2\sum_{s=0}^{\infty}e^{-s\Delta^{2}/(8\overline{\rho})}+\sum_{s=1}^{\infty}e^{-\Delta\sqrt{(t-1)/8}}\leq 2\left(1+\frac{8\overline{\rho}}{\Delta^{2}}\right)+\frac{16}{\Delta^{2}}.$ (56)
Once more, using that $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{t_{0}\log K}$ and our choice of $t_{0}$, we obtain
$\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{\log(K)}+\frac{\sqrt{8\overline{\rho}}\log(K)}{\Delta}.$
Using this in (53) together with (56) yields
$\mathbb{E}\left[\overline{R}_{\textnormal{{hdg}},T}\right]$ $\leq\left[\sqrt{\log(K)}+\frac{\sqrt{8\overline{\rho}}\log(K)}{\Delta}\right]+\left[2+\frac{16(1+\overline{\rho})}{\Delta^{2}}\right]$
$\leq 2+\frac{(1+3\overline{\rho})\log(K)}{\Delta}+\frac{16(1+\overline{\rho})}{\Delta^{2}},$
where we use $\sqrt{8}\leq 3$ for simplicity. Collecting the terms renders the required bound. ∎
Proposition B.6 (Bernstein-type regret bounds).
Let $K\geq 3$. Assume that the conditions of Lemma B.4 hold. Define $\widetilde{v}_{\max}:=\max_{k\neq k^{\star}}\widetilde{v}_{k}$.
Under the conditions of case (i) of Lemma B.4, it holds
$\mathbb{E}\big[\overline{R}_{\textnormal{{hdg}},T}\big]\leq 1+\sqrt{\log K}+\frac{4\sqrt{\frac{2}{3}}\Delta\log K+8(\widetilde{v}_{\max}+\frac{1}{3}\Delta)+16}{\Delta^{2}}.$
Under the conditions of case (ii) of Lemma B.4, and additionally, under Assumption 3(ii), it holds
$\mathbb{E}\big[\overline{R}_{\textnormal{{hdg}},T}\big]\leq 1+\sqrt{\log K}+\frac{4\sqrt{5}\overline{\rho}_{\max}\Delta\log K+16\overline{\rho}_{\max}^{2}(4\widetilde{v}_{\max}+\Delta)+16}{\Delta^{2}},$
where $\overline{\rho}_{\max}=\max_{k\neq k^{\star}}\sup_{t\geq 1}\rho^{(k)}_{t}$, with ${\rho}^{(k)}_{t}:=1+\sum_{n=1}^{t}\sqrt{\varphi_{n}\big(\big\{(\ell^{(k)}_{t}-\ell^{(k^{\star})}_{t})\big\}_{t\in\mathbb{Z}}\big)}$.
Under the conditions of case (iii) of Lemma B.4, it holds
$\mathbb{E}[\overline{R}_{\textnormal{{hdg}},T}]$ $\leq\sqrt{\widetilde{t}^{*}_{0}\log K}+\frac{16}{\Delta^{2}}+\sum_{s=\widetilde{t}^{*}_{0}}^{\infty}\exp\left(-\frac{s\widetilde{c}_{\Delta}}{2(\log s)^{\alpha}}\right),$
where $\alpha=2/\gamma_{\rm min}$,
$\widetilde{c}_{\Delta}:=\frac{\Delta^{2}}{32(\widetilde{v}_{\max}+\tfrac{1}{6}\Delta)},\qquad\widetilde{c}^{*}_{\Delta}:=\frac{\Delta^{2}}{32\,\nu^{*}},$
and $\widetilde{t}^{*}_{0}$ is defined as
$\widetilde{t}^{*}_{0}=\left\lceil\inf\left\{t\geq\max\left\{3,{e^{\alpha}},{t_{0}}\right\}\,\bigg|\,\frac{t}{(\log t)^{\alpha}}\geq\frac{2\log(2K)}{\widetilde{c}^{*}_{\Delta}}\right\}\right\rceil$
where $t_{0}$ is given in (26).
The bounds are uniform in $T$.
We proceed precisely as in the proof of Proposition B.5 using the Bernstein-type results in Corollary B.4. The proof reduces to analyzing (54) under the conditions of case (i) and case (ii) as follows.
Case (i). Define $c_{\Delta}:=\Delta^{2}/(8(\widetilde{v}_{\max}+\frac{1}{3}\Delta))$, $\nu^{*}:=\max(1,\widetilde{v}_{\max}+\frac{1}{3}\Delta)$ and set $t_{0}=\lceil\frac{8\nu^{*}\log(K)}{\Delta^{2}}\rceil$. Analogously to the proof of case (i) of Proposition B.5, we apply case (i) of Corollary B.4 and, for $t\geq t_{0}+1$, find
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$] $\leq(K-1)e^{-(t-1)c_{\Delta}}+(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/2}\,}$
$\leq\big((K-1)e^{-t_{0}c_{\Delta}}\big)e^{-(t-t_{0}-1)c_{\Delta}}+\big((K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\big)e^{-\Delta\sqrt{(t-1)\log(K)/8}}$
$\leq e^{-(t-t_{0}-1)c_{\Delta}}+e^{-\Delta\sqrt{(t-1)\log(K)/8}}.$
To obtain the last inequality we used, first, the fact that the choice of $t_{0}\geq{8\nu^{*}\log(K)}/{\Delta^{2}}\geq{\log(K)}/c_{\Delta}$ implies that $(K-1)e^{-t_{0}c_{\Delta}}\leq 1$ assuming $K\geq 3$. Second, for $t\geq t_{0}+1$ it holds that
$(K-1)\exp\left(-\Delta\sqrt{(t-1)\frac{\log K}{8}}\right)$ $\leq(K-1)\exp\left(-\Delta\sqrt{t_{0}\frac{\log K}{8}}\right)$
$\leq(K-1)\exp\left(-\sqrt{\nu^{*}(\log K)^{2}}\right)\leq 1.$
Hence,
$\sum_{t=t_{0}+1}^{\infty}\mathbb{E}[r_{t}]\leq\sum_{s=0}^{\infty}e^{-sc_{\Delta}}+\sum_{s=1}^{\infty}e^{-(\Delta/\sqrt{8})\sqrt{s}}\leq 1+\frac{1}{c_{\Delta}}+\frac{16}{\Delta^{2}},$ (57)
where we applied Lemma A.2 since $\log(K)>1$ for $K\geq 3$.
We now consider again the bound $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{t_{0}\log K}$. Let us observe that, since losses $\ell_{t}^{(k)}\in[0,1]$ for all $k\in[K]$, $v_{\max}\leq\mathbb{E}[(\ell_{t}^{(k)})^{2}]\leq 1$, and hence $\nu^{*}\leq 4/3$. Using our choice of $t_{0}$, we obtain $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{({\frac{32}{3}\log K}/{\Delta^{2}}+1)\log K}\leq\sqrt{\log K}+4\sqrt{\frac{2}{3}}\log(K)/\Delta$. Therefore, using (53) we find
$\mathbb{E}\left[\overline{R}_{\textnormal{{hdg}},T}\right]\leq\sqrt{\log K}+\frac{4\sqrt{\frac{2}{3}}\log K}{\Delta}+1+\dfrac{1}{{c_{\Delta}}}+\dfrac{16}{\Delta^{2}},$
which by definition of $c_{\Delta}$ leads to the desired bound.
Case (ii). Similarly to case (i) above, we use case (ii) of Corollary B.4 and introduce ${c}_{\Delta}:=\Delta^{2}/(16\overline{\rho}_{\max}^{2}(4\widetilde{v}_{\max}+\Delta))$ and ${\nu}^{*}:=\max(1,4\widetilde{v}_{\max}+\Delta)$. We set $t_{0}:=\left\lceil\frac{16\,\overline{\rho}_{\max}^{2}{\nu}^{*}\log K}{\Delta^{2}}\right\rceil$. For $t\geq t_{0}+1$, we thus obtain the bound
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$] $\leq\big((K-1)e^{-t_{0}{c}_{\Delta}}\big)e^{-(t-t_{0}-1){c}_{\Delta}}+\big((K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\big)e^{-\Delta\sqrt{(t-1)\log(K)/8}}$
$\leq e^{-(t-t_{0}-1){c}_{\Delta}}+e^{-\Delta\sqrt{(t-1)\log(K)/8}}.$
The first summand follows from $t_{0}\geq 16\,\overline{\rho}_{\max}^{2}{\nu}^{*}\log K/\Delta^{2}\geq\log K/{c}_{\Delta}$, while the second summand is due to noticing that
$(K-1)\exp\left(-\Delta\sqrt{\frac{16\,\overline{\rho}_{\max}^{2}{\nu}^{*}\log K}{\Delta^{2}}\cdot\frac{\log K}{8}}\right)\leq(K-1)\exp\left(-\overline{\rho}_{\max}\sqrt{2{\nu}^{*}(\log K)^{2}}\right)\leq 1,$
where we use that $\overline{\rho}_{\max}\geq 1$ by definition and $K\geq 3$.
With similar derivations are before, this time noting that ${\nu}^{*}\leq 5$ leads to $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{({80\overline{\rho}_{\max}^{2}\log K}/{\Delta^{2}}+1)\log K}$ , one finds
$\mathbb{E}\left[\overline{R}_{\textnormal{{hdg}},T}\right]\leq\sqrt{\log K}+\frac{4\sqrt{5}\overline{\rho}_{\max}\log K}{\Delta}+1+\dfrac{1}{{{c}_{\Delta}}}+\dfrac{16}{\Delta^{2}}.$
Plugging in for ${c}_{\Delta}$ and collecting terms leads to the final bound.
Case (iii). We now combine the proof strategy from previous parts with that of Proposition B.4, case (iii). To apply the bound in (47), Corollary B.4, for convenience we introduce $\alpha=2/\gamma_{\min}$. Let again $\nu^{*}:=\max(1,\widetilde{v}_{\max}+\frac{1}{6}\Delta)$, and define
$\widetilde{c}_{\Delta}:=\frac{\Delta^{2}}{32(\widetilde{v}_{\max}+\tfrac{1}{6}\Delta)},\qquad\widetilde{c}^{*}_{\Delta}:=\frac{\Delta^{2}}{32\,\nu^{*}},$
as well as
$\widetilde{t}^{*}_{0}=\left\lceil\inf\left\{t\geq\max\left\{3,{e^{\alpha}},{t_{0}}\right\}\,\bigg|\,\frac{t}{(\log t)^{\alpha}}\geq\frac{2\log(2K)}{\widetilde{c}^{*}_{\Delta}}\right\}\right\rceil,$
which is well-defined thanks to Lemma A.3. For $t\geq\widetilde{t}^{*}_{0}+1$, we obtain
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right$] $\leq 2(K-1)\exp\left(-\frac{(t-1)\widetilde{c}_{\Delta}}{(\log(t-1))^{\alpha}}\right)+(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/2}}$
$=I+II,$
where we can study terms $I$ and $II$ separately. With regards to the former, note that
$I$ $\leq\exp\left(\log(2K)-\frac{(t-1)\widetilde{c}_{\Delta}}{(\log(t-1))^{\alpha}}\right)$
$\leq\exp\left(\frac{(t-1)\widetilde{c}^{*}_{\Delta}}{2(\log(t-1))^{\alpha}}-\frac{(t-1)\widetilde{c}_{\Delta}}{(\log(t-1))^{\alpha}}\right)\leq\exp\left(-\frac{(t-1)\widetilde{c}_{\Delta}}{2(\log(t-1))^{\alpha}}\right),$
where we have used that $\widetilde{c}^{*}_{\Delta}\leq\widetilde{c}_{\Delta}$ by definition of $\nu^{*}$. Turning to $II$, much like in previous parts, we can split the exponential factor to have
$II=\big((K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}\big)e^{-\Delta\sqrt{(t-1)\log(K)/8}}.$
We now use that, since $\widetilde{t}^{*}_{0}\geq 3$, it holds
$\widetilde{t}^{*}_{0}\geq(\log\widetilde{t}^{*}_{0})^{\alpha}\log(2K)\frac{64\nu^{*}}{\Delta^{2}}\geq\log(2K)\frac{64\nu^{*}}{\Delta^{2}},$
and hence
$(K-1)e^{-\Delta\sqrt{(t-1)\log(K)/8}}$ $\leq(K-1)\exp\left(-\Delta\sqrt{\log(2K)\frac{64\nu^{*}}{\Delta^{2}}\cdot\frac{\log K}{8}}\right)\leq 1$
when $t\geq\widetilde{t}^{*}_{0}+1$, since $\log(2K)\geq\log K$. Putting the two bounds together, one finds, for $t$ sufficiently large,
$\sum_{k\neq k^{\star}}\mathbb{E}\left[\omega^{(k)}_{\textnormal{{hdg}},t}\right]\leq\exp\left(-\frac{(t-1)\widetilde{c}_{\Delta}}{2(\log(t-1))^{\alpha}}\right)+\exp\left(-\Delta\sqrt{(t-1)\frac{\log K}{8}}\right).$ (58)
Once more, we use the result that $\overline{R}_{\textnormal{{hdg}},t_{0}}\leq\sqrt{\widetilde{t}^{*}_{0}\log K}$ to control the warm-up period and Lemma A.2. We obtain an overall regret bound
$\sqrt{\widetilde{t}^{*}_{0}\log K}+\frac{16}{\Delta^{2}}+\sum_{s=\widetilde{t}^{*}_{0}}^{\infty}\exp\left(-\frac{s\widetilde{c}_{\Delta}}{2(\log s))^{\alpha}}\right).$
B.4.1 Combined Bounds for Hedge
We can now obtain Theorem 3.2 by incorporating the bounds previously derived.
Part (i) follows immediately by combining Proposition B.5(i) and Proposition B.6(i). Similarly, part (ii) is obtained by combining Proposition B.5(ii) and Proposition B.6(ii), with the relevant constants and $t_{0}$ defined in the latter. ∎
Appendix C Doubling-trick Hedge
Constant Hedge can be modified to work in setups where the time horizon $T$ is not known in advance, or when the learning process is intended to continue, by using the so-called doubling trick (Cesa-Bianchi and Lugosi,, 2006; Shalev-Shwartz,, 2012). The idea is to partition time into phases of exponentially increasing length, and to restart Hedge at the beginning of each phase with a learning rate tuned for that phase length.
Specifically, let phase $r$ cover rounds $t\in[2^{r-1},2^{r}-1]$, and set
$\eta_{r}=\sqrt{\frac{8\log K}{S_{T}^{2}2^{r-1}}}.$
At the start of phase $r$, the forecaster resets the cumulative losses to zero and applies Hedge with constant rate $\eta_{r}$ until the end of the phase. Each phase therefore behaves as if the horizon were $2^{r-1}$ and the total number of phases up to round $T$ is at most $\lfloor\log_{2}T\rfloor+1$. Summing the regret bounds in (7) over all completed phases yields $\overline{R}_{T}\leq S_{T}\sqrt{2T\log K}+\mathcal{O}(S_{T}\log K)$, which matches the $\mathcal{O}(\sqrt{T\log K})$ worst-case rate of Hedge with known $T$.
The doubling trick thus provides a simple, parameter-free procedure for adaptively restarting Hedge as time progresses, while maintaining the same asymptotic performance guarantees as the optimal constant-rate scheme with a known horizon.
Appendix D Direct Combination Methods
We focus on linear weighting schemes, known to be parsimonious and interpretable. We consider two “direct” approaches, the so-called Simple Averaging (SA) and the rolling Mean Squared Error (rollMSE) methods, which are readily implementable online and serve as benchmarks in our empirical experiments in Section 5.
Simple Averaging (SA).
One can assign equal time-invariant weights to all $K$ experts
$\omega^{(k)}_{\mathrm{SA},t}=\frac{1}{K}\quad\text{for all }k\in[K],\kern 5.0ptt\in[T].$ (59)
SA is straightforward and can reduce variance when many experts are accurate. However, it is inflexible, as it does not adapt to variability in time in relative performance across experts and implicitly relies on a subset of experts performing well at all times. Moreover, it also requires that no expert faces catastrophically high losses over the forecasting window, as this can increase the forecaster’s losses unless additional tools are used (for example, trimming or winsorization).
Rolling MSE (rollMSE).
The rolling MSE scheme is specifically tailored to the setting of squared losses for experts. At decision time $t\in[T]$ and forecasting horizon $h\in[H]$, only losses from decisions $\tau\leq t-h$ have realized (their outcomes are revealed at $\tau+h\leq t$). Fix a window length $r\geq 1$ and define the effective window length
$r(t,h):=\min\{r,(t-h)_{+}\},\quad(x)_{+}:=\max\{x,0\}.$
If $r(t,h)=0$ ($t\leq h$), use an initialization rule (for example, equal weights). Otherwise, for each expert $k\in[K]$ set
$\mathrm{MSE}^{(k)}_{t,h,r}=\frac{1}{r(t,h)}\sum_{\tau=t-h-r(t,h)+1}^{t-h}\ell^{(k)}_{\tau,h}=\frac{1}{r(t,h)}\sum_{\tau=t-h-r(t,h)+1}^{t-h}\left(Y_{\tau+h}-\widehat{Y}^{(k)}_{\tau+h}\right)^{2},$
and construct the rolling MSE weights according to
$\omega^{(k)}_{\mathrm{rollMSE},t}=\frac{\left({\mathrm{MSE}}^{(k)}_{t,h}+\varepsilon\right)^{-1}}{\sum_{j=1}^{K}\left({\mathrm{MSE}}^{(j)}_{t,h}+\varepsilon\right)^{-1}},$
with $\varepsilon>0$ a small constant.
This approach has proven effective in practice, and does not require estimating forecast error correlations (Timmermann,, 2006). The key practical issue is the selection of the window size $r$. Selecting a window size too large can produce weights that react too slowly to sudden changes in loss values across experts, while a small $r$ can overreact to recent performance. Moreover, the weighting of MSEs within the window is imposed to be uniform. For the extreme case $r=1$ ($t>h$ and $r(t,h)=1$), the weights depend only on the most recent realized $h$-ahead error, that is
$\omega^{(k)}_{\mathrm{rollMSE},t}$ $=\dfrac{\left(\ell^{(k)}_{t-h}+\varepsilon\right)^{-1}}{\sum_{j=1}^{K}\left(\ell^{(j)}_{t-h}+\varepsilon\right)^{-1}}$
$=\dfrac{\left(\left(Y_{(t-h)+h}-\widehat{Y}^{(k)}_{(t-h)+h}\right)^{2}+\varepsilon\right)^{-1}}{\sum_{j=1}^{K}\left(\left(Y_{(t-h)+h}-\widehat{Y}^{(j)}_{(t-h)+h}\right)^{2}+\varepsilon\right)^{-1}}=\dfrac{\left(\left(Y_{t}-\widehat{Y}^{(k)}_{t}\right)^{2}+\varepsilon\right)^{-1}}{\sum_{j=1}^{K}\left(\left(Y_{t}-\widehat{Y}^{(j)}_{t}\right)^{2}+\varepsilon\right)^{-1}}.$
Appendix E Algorithms
In this section, we provide pseudo-codes for the algorithm implementation of the ensemble combination schemes discussed in Section 3 and then used in Section 5. As we only evalute one-step-ahead forecasts ($h=1$), we simplify the notation outlined in Section 2 as follows:
$\omega^{(1)}_{t+1}\equiv\omega^{(k)}_{t,1},\quad\bm{\omega}_{t+1}\equiv\bm{\omega}_{t,1},\quad\ell^{(k)}_{t+1}\equiv\ell^{(k)}_{t,1},\quad L^{(k)}_{t+1}\equiv L^{(k)}_{t,1}$.
An explicit subscript for each combination method is also added (see the notation of Section 3).
Algorithm 1 – Follow-the-Leader (FTL)
Ensemble of experts indexed by $k\in[K]$
Initialize cumulative losses as $L^{(k)}_{1}=0$ for all $k\in[K]$
Initialize weights as $w_{\textnormal{{ftl}},2}^{(k)}=1/K$ for all $k\in[K]$ $\triangleright$ Equal weighting init
for $t=1,2,\ldots,T$ do
   //˜ Prediction step  ˜
   Experts make predictions $\{\widehat{Y}^{(k)}_{t+1}\}_{k=1}^{K}$
   Ensemble prediction is $\widehat{Y}_{t+1}:=\sum_{k=1}^{K}\omega_{\textnormal{{ftl}},t+1}^{(k)}\widehat{Y}_{t+1}^{(k)}$
   Outcome $Y_{t+1}$ is realized
   //˜ Weights update step  ˜
   Each $k$th expert, $k\in[K]$, updates its cumulative loss $L^{(k)}_{t+1}=L^{(k)}_{t}+\ell(Y_{t+1},\widehat{Y}_{t+1}^{(k)})$
   Compute set of experts with minimal cumulative loss, $\textnormal{{ftl}}_{t+1}=\arg\min_{k\in[K]}L^{(k)}_{t+1}$
   Set $\omega_{\textnormal{{ftl}},t+2}^{(k)}=\mathds{1}\{k\in\textnormal{{ftl}}_{t+1}\}/\lvert\textnormal{{ftl}}_{t+1}\rvert$ for all $k\in[K]$
output Ensemble predictions $\{\widehat{Y}_{t+1}\}_{t=1}^{T}$, ensemble weights $\{\bm{\omega}_{\textnormal{{ftl}},{t+1}}\}_{t=1}^{T}$
Algorithm 2 – Constant Hedge
An ensemble of experts indexed by $k\in[K]$, constant learning rate $\eta>0$
Initialize weights as $w_{\textnormal{{hdg}},2}^{(k)}=1/K$ for all $k\in[K]$ $\triangleright$ Equal weighting init
for $t=1,2,\ldots,T$ do
   //˜ Prediction step  ˜
   Experts make predictions $\{\widehat{Y}^{(k)}_{t+1}\}_{k=1}^{K}$
   Ensemble prediction is $\widehat{Y}_{t+1}:=\sum_{k=1}^{K}\omega_{\textnormal{{hdg}},t+1}^{(k)}\widehat{Y}_{t+1}^{(k)}$
   Outcome $Y_{t+1}$ is realized
   //˜ Weights update step  ˜
   Each $k$th expert, $k\in[K]$, faces instantaneous loss $\ell^{(k)}_{t+1}=\ell(Y_{t+1},\widehat{Y}_{t+1}^{(k)})$
   Compute $v^{(k)}_{t+1}:=\omega^{(k)}_{\textnormal{{hdg}},t+1}\exp(-\eta\ell^{(k)}_{t+1})$ for all $k\in[K]$ $\triangleright$ Pre-normalization weights
   Set $\omega_{\textnormal{{hdg}},t+2}^{(k)}=v^{(k)}_{t+1}/\sum_{k=1}^{K}v^{(k)}_{t+1}$ for all $k\in[K]$
output Ensemble predictions $\{\widehat{Y}_{t+1}\}_{t=1}^{T}$, ensemble weights $\{\bm{\omega}_{\textnormal{{hdg}},t+1}\}_{t=1}^{T}$
Note: Online weight updates using instantaneous losses $\ell^{(k)}_{t+1}$ are more numerically stable compared to calculating Hedge weights based on cumulative losses.
Algorithm 3 – Decreasing Hedge
An ensemble of experts indexed by $k\in[K]$, learning rate scale $c_{0}>0$ (for example $c_{0}=2$, based on worst-case bounds in Mourtada and Gaïffas,, 2019)
Initialize weights as $w_{\textnormal{{dh}},2}^{(k)}=1/K$ for all $k\in[K]$ $\triangleright$ Equal weighting init
for $t=1,2,\ldots,T$ do
   //˜ Prediction step  ˜
   Experts make predictions $\{\widehat{Y}^{(k)}_{t+1}\}_{k=1}^{K}$
   Ensemble prediction is $\widehat{Y}_{t+1}:=\sum_{k=1}^{K}\omega_{\textnormal{{dh}},t+1}^{(k)}\widehat{Y}_{t+1}^{(k)}$
   Outcome $Y_{t+1}$ is realized
   //˜ Weights update step  ˜
   Each $k$th expert, $k\in[K]$, faces instantaneous loss $\ell^{(k)}_{t+1}=\ell(Y_{t+1},\widehat{Y}_{t+1}^{(k)})$
   Each $k$th expert, $k\in[K]$, updates its cumulative loss $L^{(k)}_{t+1}=L^{(k)}_{t}+\ell(Y_{t+1},\widehat{Y}_{t+1}^{(k)})$
   Find $L^{(k^{\star})}_{t+1}=\min_{k\in[K]}L^{(k)}_{t+1}$
   Set $\eta_{t+1}=c_{0}\sqrt{\log(K)/(t+1)}$ $\triangleright$ Decreasing learning rate
   Compute $v^{(k)}_{t+1}:=\exp(-\eta_{t+1}(L^{(k)}_{t+1}-L^{(k^{\star})}_{t+1}))$ for all $k\in[K]$ $\triangleright$ Pre-normalization weights
   Set $\omega_{\textnormal{{dh}},t+2}^{(k)}=v^{(k)}_{t+1}/\sum_{k=1}^{K}v^{(k)}_{t+1}$ for all $k\in[K]$
output Ensemble predictions $\{\widehat{Y}_{t+1}\}_{t=1}^{T}$, ensemble weights $\{\bm{\omega}_{\textnormal{{dh}},t+1}\}_{t=1}^{T}$
Note: Hedge weights are calculated by centering with the minimum expert cumulative loss $L^{(k^{\star})}_{t+1}$ to avoid numerical issues.
Algorithm 4 – AdaHedge
An ensemble of experts indexed by $k\in[K]$
Initialize cumulative losses as $L^{(k)}_{1}=0$ for all $k\in[K]$
Initialize cumulative mixability gab $\nabla_{1}=0$ and ${M}_{1}=0$
for $t=1,2,\ldots,T$ do
   //˜ Weight computation  ˜
   Find $L^{(k^{\star})}_{t}=\min_{k\in[K]}L^{(k)}_{t}$
   if $\nabla_{t}=0$ then $\triangleright$ Pre-normalization weights
    Set $\eta_{t}=\infty$ and $v^{(k)}_{t}=\mathds{1}\{L^{(k)}_{t}=L^{(k^{\star})}_{t}\}$
   else
    Set $\eta_{t}=\log(K)/\nabla_{t}$ and $v^{(k)}_{t}=\exp\left(-\eta_{t}(L^{(k)}_{t}-L^{(k^{\star})}_{t})\right)$ for all $k\in[K]$
   Compute $\overline{v}_{t}=\sum_{k=1}^{K}v^{(k)}_{t}$
   Set $\omega_{\textnormal{{ah}},t+1}^{(k)}=v^{(k)}_{t}/\overline{v}_{t}$ for all $k\in[K]$ $\triangleright$ Current period weights
   //˜ Prediction step  ˜
   Experts make predictions $\{\widehat{Y}^{(k)}_{t+1}\}_{k=1}^{K}$
   Ensemble prediction is $\widehat{Y}_{t+1}:=\sum_{k=1}^{K}\omega_{\textnormal{{ah}},t+1}^{(k)}\widehat{Y}_{t+1}^{(k)}$
   Outcome $Y_{t+1}$ is realized
   //˜ Mixability gap update step  ˜
   Each $k$th expert, $k\in[K]$, faces instantaneous loss $\ell^{(k)}_{t+1}=\ell(Y_{t+1},\widehat{Y}_{t+1}^{(k)})$
   Each $k$th expert, $k\in[K]$, updates its cumulative loss $L^{(k)}_{t+1}=L^{(k)}_{t}+\ell^{(k)}_{t+1}$
   Find $L^{(k^{\star})}_{t+1}=\min_{k\in[K]}L^{(k)}_{t+1}$
   Compute forecaster’s loss $\overline{\ell}_{t+1}=\bm{\omega}_{\textnormal{{AH}},t+1}^{\top}\bm{\ell}_{t+1}$
   Compute $M_{t+1}=L^{(k^{\star})}_{t+1}-\eta_{t}^{-1}\log(\overline{v}_{t+1}/K)$ $\triangleright$ Numerically stable mixing
   Update mixability gap $\nabla_{t+1}=\nabla_{t}+\max(0,\overline{\ell}_{t+1}-(M_{t+1}-M_{t}))$
output Ensemble predictions $\{\widehat{Y}_{t+1}\}_{t=1}^{T}$, ensemble weights $\{\bm{\omega}_{\textnormal{{ah}},t+1}\}_{t=1}^{T}$
Note: The update of the mixability gap includes a $\max(0,\cdot)$ operation to avoid numerical violations of Jensen’s inequality. The mixability gap update used is based on the upper bound from Lemma 2 in de Rooij et al., (2014).
Appendix F Summary of Data
The dataset is collected from various sources, including the Federal Reserve Bank of St. Louis Monthly (FRED-MD) and Quarterly (FRED-QD) Macroeconomic Databases (see Stock and Watson,, 1996, 2002 for more details), as well as daily series sourced from Refinitiv Datastream. The selection of predictors in the medium-MD dataset includes indicators from ten macroeconomic and financial categories (see Stock and Watson,, 1996, 2002). The preprocessing of macroeconomic indicators follows the standard guidelines outlined in McCracken and Ng, (2016, 2020). The full data sample spans a period from January 1st, 1990, to December 31st, 2019. We concentrate on the evaluation up until the Great Financial Crisis, covering the period from 1990Q1 to 2007Q4. The testing and online learning sample spans 2008Q1-2019Q4, encompassing a total of 48 forecasting rounds.
Start Date T Code Name Description
Quarterly
31/03/1959 5 GDPC1 Y Real Gross Domestic Produce
Monthly
30/01/1959 5 INDPRO XM1 Industrial Production Index
30/01/1959 5 PAYEMS XM4 Payroll All Employees: Total nonfarm
30/01/1959 4 HOUST XM5 Housing Starts: Total New Privately Owned
30/01/1959 5 RETAILx XM7 Retail and Food Services Sales
31/01/1973 5 TWEXMMTH XM11 Nominal effective exchange rate US
30/01/1959 2 FEDFUNDS XM12 Effective Federal Funds Rate
30/01/1959 1 BAAFFM XM14 Moody’s Baa Corporate Bond Minus FEDFUNDS
30/01/1959 1 COMPAPFFx XM15 3-Month Commercial Paper Minus FEDFUNDS
30/01/1959 2 CUMFNS XM2 Capacity Utilization: Manufacturing
30/01/1959 2 UNRATE XM3 Civilian Unemployment Rate
30/01/1959 5 DPCERA3M086SBEA XM6 Real personal consumption expenditures
30/01/1959 5 AMDMNOx XM8 New Orders for Durable Goods
31/01/1978 2 UMCSENTx XM9 Consumer Sentiment Index
30/01/1959 6 WPSFD49207 XM10 PPI: Finished Goods
30/01/1959 1 AAAFFM XM13 Moody’s Aaa Corporate Bond Minus FEDFUNDS
30/01/1959 1 TB3SMFFM XM16 3-Month Treasury C Minus FEDFUNDS
30/01/1959 1 T10YFFM XM17 10-Year Treasury C Minus FEDFUNDS
30/01/1959 2 GS1 XM18 1-Year Treasury Rate
30/01/1959 2 GS10 XM19 10-Year Treasury Rate
30/01/1959 1 GS10-TB3MS XM20 10-Year Treasury Rate - 3-Month Treasury Bill
Daily
30/01/1959 8 DJINDUS XD3 DJ Industrial price index
31/12/1963 8 S&PCOMP XD1 S&P500 price index
01/05/1982 1 ISPCS00-S&PCOMP* XD2 S&P500 basis spread
11/09/1989 8 SP5EIND XD4 S&P Industrial price index
31/12/1969 8 GSCITOT XD5 Spot commodity price index
10/01/1983 8 CRUDOIL XD6 Spot price oil
02/01/1979 8 GOLDHAR XD7 Spot price gold
30/03/1982 8 WHEATSF XD8 Spot price wheat
01/11/1983 8 COCOAIC,COCINU** XD9 Spot price cocoa
30/03/1983 1 NCLC.03-NCLC.01 XD10 Futures price oil term structure
30/10/1978 1 NGCC.03-NGCC.01 XD11 Futures price gold term structure
02/01/1975 1 CWFC.03-CWFC.01 XD12 Futures price wheat term structure
02/01/1973 1 NCCC.03-NCCC.01 XD13 Futures price cocoa term structure
Notes: “Start Date” is the date for which the series is first available (before data transformations). Following McCracken and Ng, (2016) and McCracken and Ng, (2020), the transformation codes in column “T” indicate with D for difference and log for natural logarithm, 1: none, 2: D, 3: DD, 4: Log, 5: Dlog, 6: DDlog, 7: percentage change, 8: GARCH volatility. “Code” is the code in the FRED-QD and FRED-MD datasets for quarterly and monthly data, and the Datastream mnemonic for the remaining frequencies. Missing values due to public holidays are interpolated by averaging over the previous five observations. *: Available until 20/09/2021. **: Average before 29/12/2017, COCINUS mean adjusted thereafter.
Table F.1: Input and output variables, frequencies, and transformations (adapted from Ballarin et al., 2024a ).
Appendix G Summary of Models
Model Name Description Specification
Mean Unconditional mean of outcome series over the estimation sample.
  AR(1) Autoregressive model of the output series estimated using OLS.
  DFM A Stock aggregation, VAR(1) factor process. Factors: 10
  DFM B Almon aggregation, VAR(1) factor process. Same as DFM A
  S-MFESN A S-MFESN model: Sparse-normal $\widetilde{A}$, sparse-uniform $\widetilde{C}$, $\widetilde{\bm{\zeta}}=0$. Isotropic ridge regression fit. Reservoir dimension: 30 Sparsity: 33.3% $\rho=0.5$, $\gamma=1$, $\alpha=0.1$
  S-MFESN B Same as S-MFESN A except for larger reservoir dimension and lower sparsity ratio. Reservoir dimension: 120 Sparsity: 8.3% $\rho=0.5$, $\gamma=1$, $\alpha=0.1$
[…]
EN-MFESN-RP Applied to all types of MFESN: 1000 distinct models generated with independently randomly drawn reservoir state coefficients. Same specification as the corresponding baseline MFESN model.
[…]
Table G.1: Models and ensembles applied in empirical forecasting exercises (for non-ensemble models see Table 4.1 in Ballarin et al., 2024a ). MFESN hyperparameters are defined for normalized state equations (11)-(12).
G.1 Combination Weights Dynamics
[…]
Our findings indicate that weight dynamics can differ significantly across different classes of MFESNs ensembles. In the upper row of Figure G.1, both AdaHedge and FTL weights for the S-MFESN A ensembles change substantially over the forecasting exercise. For AdaHedge (panel (a) in Figure G.1) the weight vector $\bm{\omega}_{\textnormal{{hdg}},t}$ remains very close to its equal-weight initialization until the second quarter of 2010. After that, a clearer ranking starts to emerge slowly, but ceases to stabilize. A similar pattern is visible for the Follow-the-Leader (panel (b) of Figure G.1), whose weights $\bm{\omega}_{\textnormal{{FTL}},t}$ fluctuate heavily until 2012Q2, after which the leader remains unchanged. By contrast, the dynamics of M-MFESN A ensembles’ weights are more stable. Panels (c) and (d) of Figure G.1 demonstrate that AdaHedge and FTL weights settle quickly and change slowly over time. For AdaHedge, the weights $\bm{\omega}_{\textnormal{{hdg}},t}$ stabilize after the end of the Great Financial Crisis downturn. Over 80% of the cumulative weight concentrates on just 5 experts, and the top-ranked expert receives over 30% of the weight as early as 2009Q1. Furthermore, the weights $\bm{\omega}_{\textnormal{{FTL}},t}$ of the FTL demonstrate quick convergence to a stable leader within less than one year from the start of the prediction exercise, signaling that the ensemble contains an M-MFESN A model which achieves consistent optimality in terms of MSFE.
It is important to emphasize that these marked differences in weight dynamics do not imply large differences in overall forecasting performance. Table 2 reports that MFESN A-type results for FTL and AdaHedge across single- and multi-reservoir architectures show much smaller MSFE gaps than, for example, across A- and B-type architectures. From an interpretability perspective, more stable weights provide clearer insight into the mechanism of the ensemble: Figure G.1(d) strongly suggests an online form of model selection carried out iteratively during the forecasting exercise. An open question is hence whether there exist inherent trade-offs between ensemble “interpretability” (e.g., provably simple weights dynamics for model instances from given classes of nonlinear models) and predictive performance. We defer the treatment of this question to future work.
(a) S-MFESN A – AdaHedge
(b) S-MFESN A – Follow-the-Leader
(c) M-MFESN A – AdaHedge
(d) M-MFESN A – Follow-the-Leader
[…]
(a) S-MFESN B – AdaHedge
(b) S-MFESN B – Follow-the-Leader
(c) M-MFESN B – AdaHedge
(d) M-MFESN B – Follow-the-Leader
[…]
Appendix H Additional Tables
In this appendix, we provide results for the additional empirical experiments conducted under multistep forecasting scenarios. Tables J.1-J.3 report the relative MSE of the quarterly U.S. GDP growth predictions with respect to the sample mean for the exercise with the forecasting horizons taken as $h=2$, $h=4$, and $h=8$ quarters ahead, respectively. All competing models were implemented to produce the iterative $h$-steps-ahead forecasts (see Ballarin et al., 2024a for more details regarding the ESN-based models). It is important to emphasize that the hyperparameters for all the baseline MFESN models are tuned with cross-validation only for the prediction at $h=1$. This implies that baseline models are not expected to produce their best results with respect to the unconditional mean, AR(1), and DFM benchmarks. Our goal is to inspect the changes in the accuracy of the combinations of expert forecasts with respect to the performance offered by the standalone baseline models. We emphasize that in all combination schemes, the weights of the expert models in an ensemble are based on the past behavior in the $h$ steps ahead forecasting exercise. Our findings highlight that there are meaningful changes in relative improvements resulting from the use of combination strategies as the forecast horizon increases.
More specifically, for $h=2$, similar to the case of $h=1$ in Table 2, the FTL combination strategy together with the AdaHedge algorithm again appear to dominate all other combination schemes for the corresponding baseline models, significantly improving the forecasting accuracy of the latter once used in an ensemble. Starting from $h=4$ onward, several differences are noticeable. First, the expert selected by FTL as the leader (that is, the one that produces the most accurate predictions measured in terms of the cumulative loss) ceases to be the best predictor in the subsequent period. While FTL can adapt quickly to changes in the best expert for one-step-ahead predictions, its ability to track the best expert over longer horizons diminishes. Second, the AdaHedge and DecHedge strategies maintain strong performance for longer prediction horizons, especially for the case of EN-MFESN-RP ensembles, where only the parameters of the state equation are resampled.
We conclude by noting that further refinements could be potentially explored for $h>1$: average losses for all steps up to $h$ could be used in the weights update rule; discounted cumulative losses up to step $h$ may be employed; hyperparameters of the baseline models could be subject to cross-validation, and others. We leave these avenues for potential improvement for future work.
—  Horizon $h=2$  —
[…]
Table H.1: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=4$  —
[…]
Table H.2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=8$  —
[…]
Table H.3: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
Start Date T Code Name Description
Quarterly
31/03/1959 5 GDPC1 Y Real Gross Domestic Produce
Monthly
30/01/1959 5 INDPRO XM1 Industrial Production Index
30/01/1959 5 PAYEMS XM4 Payroll All Employees: Total nonfarm
30/01/1959 4 HOUST XM5 Housing Starts: Total New Privately Owned
30/01/1959 5 RETAILx XM7 Retail and Food Services Sales
31/01/1973 5 TWEXMMTH XM11 Nominal effective exchange rate US
30/01/1959 2 FEDFUNDS XM12 Effective Federal Funds Rate
30/01/1959 1 BAAFFM XM14 Moody’s Baa Corporate Bond Minus FEDFUNDS
30/01/1959 1 COMPAPFFx XM15 3-Month Commercial Paper Minus FEDFUNDS
30/01/1959 2 CUMFNS XM2 Capacity Utilization: Manufacturing
30/01/1959 2 UNRATE XM3 Civilian Unemployment Rate
30/01/1959 5 DPCERA3M086SBEA XM6 Real personal consumption expenditures
30/01/1959 5 AMDMNOx XM8 New Orders for Durable Goods
31/01/1978 2 UMCSENTx XM9 Consumer Sentiment Index
30/01/1959 6 WPSFD49207 XM10 PPI: Finished Goods
30/01/1959 1 AAAFFM XM13 Moody’s Aaa Corporate Bond Minus FEDFUNDS
30/01/1959 1 TB3SMFFM XM16 3-Month Treasury C Minus FEDFUNDS
30/01/1959 1 T10YFFM XM17 10-Year Treasury C Minus FEDFUNDS
30/01/1959 2 GS1 XM18 1-Year Treasury Rate
30/01/1959 2 GS10 XM19 10-Year Treasury Rate
30/01/1959 1 GS10-TB3MS XM20 10-Year Treasury Rate - 3-Month Treasury Bill
Daily
30/01/1959 8 DJINDUS XD3 DJ Industrial price index
31/12/1963 8 S&PCOMP XD1 S&P500 price index
01/05/1982 1 ISPCS00-S&PCOMP* XD2 S&P500 basis spread
11/09/1989 8 SP5EIND XD4 S&P Industrial price index
31/12/1969 8 GSCITOT XD5 Spot commodity price index
10/01/1983 8 CRUDOIL XD6 Spot price oil
02/01/1979 8 GOLDHAR XD7 Spot price gold
30/03/1982 8 WHEATSF XD8 Spot price wheat
01/11/1983 8 COCOAIC,COCINU** XD9 Spot price cocoa
30/03/1983 1 NCLC.03-NCLC.01 XD10 Futures price oil term structure
30/10/1978 1 NGCC.03-NGCC.01 XD11 Futures price gold term structure
02/01/1975 1 CWFC.03-CWFC.01 XD12 Futures price wheat term structure
02/01/1973 1 NCCC.03-NCCC.01 XD13 Futures price cocoa term structure
Notes: “Start Date” is the date for which the series is first available (before data transformations). Following McCracken and Ng, (2016) and McCracken and Ng, (2020), the transformation codes in column “T” indicate with D for difference and log for natural logarithm, 1: none, 2: D, 3: DD, 4: Log, 5: Dlog, 6: DDlog, 7: percentage change, 8: GARCH volatility. “Code” is the code in the FRED-QD and FRED-MD datasets for quarterly and monthly data, and the Datastream mnemonic for the remaining frequencies. Missing values due to public holidays are interpolated by averaging over the previous five observations. *: Available until 20/09/2021. **: Average before 29/12/2017, COCINUS mean adjusted thereafter.
Table H.4: Input and output variables, frequencies, and transformations (adapted from Ballarin et al., 2024a ).
Appendix I Summary of Models
Model Name Description Specification
Mean Unconditional mean of outcome series over the estimation sample.
  AR(1) Autoregressive model of the output series estimated using OLS.
  DFM A Stock aggregation, VAR(1) factor process. Factors: 10
  DFM B Almon aggregation, VAR(1) factor process. Same as DFM A
  S-MFESN A S-MFESN model: Sparse-normal $\widetilde{A}$, sparse-uniform $\widetilde{C}$, $\widetilde{\bm{\zeta}}=0$. Isotropic ridge regression fit. Reservoir dimension: 30 Sparsity: 33.3% $\rho=0.5$, $\gamma=1$, $\alpha=0.1$
  S-MFESN B Same as S-MFESN A except for larger reservoir dimension and lower sparsity ratio. Reservoir dimension: 120 Sparsity: 8.3% $\rho=0.5$, $\gamma=1$, $\alpha=0.1$
[…]
EN-RP-ESN Applied to all types of MFESN: 1000 distinct models generated with independently randomly drawn reservoir state coefficients. Same specification as the corresponding baseline MFESN model.
[…]
Table I.1: Models and ensembles applied in empirical forecasting exercises (for non-ensemble models see Table 4.1 in Ballarin et al., 2024a ). MFESN hyperparameters are defined for normalized state equations (11)-(12).
Appendix J Additional Tables
In this appendix, we provide results for the additional empirical experiments conducted under multistep forecasting scenarios. Tables J.1-J.3 report the relative MSE of the quarterly U.S. GDP growth predictions with respect to the sample mean for the exercise with the forecasting horizons taken as $h=2$, $h=4$, and $h=8$ quarters ahead, respectively. All competing models were implemented to produce the iterative $h$-steps-ahead forecasts (see Ballarin et al., 2024a for more details regarding the ESN-based models). It is important to emphasize that the hyperparameters for all the baseline MFESN models are tuned with cross-validation only for the prediction at $h=1$. This implies that baseline models are not expected to produce their best results with respect to the unconditional mean, AR(1), and DFM benchmarks. Our goal is to inspect the changes in the accuracy of the combinations of expert forecasts with respect to the performance offered by the standalone baseline models. We emphasize that in all combination schemes, the weights of the expert models in an ensemble are based on the past behavior in the $h$ steps ahead forecasting exercise. Our findings highlight that there are meaningful changes in relative improvements resulting from the use of combination strategies as the forecast horizon increases.
More specifically, for $h=2$, similar to the case of $h=1$ in Table 2, the FTL combination strategy together with the AdaHedge algorithm again appear to dominate all other combination schemes for the corresponding baseline models, significantly improving the forecasting accuracy of the latter once used in an ensemble. Starting from $h=4$ onward, several differences are noticeable. First, the expert selected by FTL as the leader (that is, the one that produces the most accurate predictions measured in terms of the cumulative loss) ceases to be the best predictor in the subsequent period. While FTL can adapt quickly to changes in the best expert for one-step-ahead predictions, its ability to track the best expert over longer horizons diminishes. Second, the AdaHedge and DecHedge strategies maintain strong performance for longer prediction horizons, especially for the case of EN-RP-ESN ensembles, where only the parameters of the state equation are resampled.
We conclude by noting that further refinements could be potentially explored for $h>1$: average losses for all steps up to $h$ could be used in the weights update rule; discounted cumulative losses up to step $h$ may be employed; hyperparameters of the baseline models could be subject to cross-validation, and others. We leave these avenues for potential improvement for future work.
—  Horizon $h=2$  —
[…]
Table J.1: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=4$  —
[…]
Table J.2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=8$  —
[…]
Table J.3: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.